### v1

Abstract
Context. Together with gravity, turbulence, and stellar feedback, magnetic fields (B-fields) are thought to play a critical role in the evolution of molecular clouds and star formation processes. Polarization of thermal dust emission is a popular tracer of B-fields in star-forming regions.
Aims. We aim to map the morphology and measure the strength of B-fields of the nearby molecular cloud, rho Ophiuchus-A ( Oph-A), and then to understand the role of B-fields in regulating star formation and shaping the cloud.
Methods. We have analyzed the far-infrared (FIR) polarization of thermal dust emission observed by SOFIA/HAWC at 89 and 154 m toward the densest part of Oph-A, which is irradiated by the nearby B3/4 star, Oph-S1. These FIR polarimetric maps cover an area of (corresponding to pc2) with an angular resolution of 7.8 and 13.6, respectively.
Results. The Oph-A cloud exhibits well-ordered B-fields with magnetic orientations mainly perpendicular to the ridge of the cloud toward the densest region. We obtain a map of B-field strengths in the range of 0.2-2.5 mG, using the Davis–Chandrasekhar-Fermi method. The B-fields are strongest at the densest part of the cloud, which is associated with the starless core SM1, and decreases toward the outskirts of the cloud. By calculating the map of the mass-to-flux ratio, Alfvén Mach number, and plasma parameter in Oph-A, we find that the cloud is predominantly magnetically sub-critical, sub-Alfvénic, which implies that the cloud is supported by strong B-fields that dominate over gravity, turbulence, and thermal gas energy. Measured B-field strengths at two densest subsregions using other methods that account for the compressible mode are relatively lower than that measured with the DCF method but do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. A virial analysis suggests that the cloud is gravitationally unbound, which is consistent with the previous detection of numerous starless cores in the cloud. Comparing the magnetic pressure with the radiation pressure from the Oph-S1 star, we find that B-fields are sufficiently strong to support the cloud against radiative feedback and to regulate the shape of the cloud.
Key Words.:
ISM: magnetic fields – star: formation - ISM: clouds – ISM: individual objects: Rho Oph-A1 Introduction
Stars like our Sun form in the cold and dense parts of molecular clouds, which consist of filaments, clumps, and dense cores (Shu et al., 1987; Lada & Lada, 2003; Kennicutt & Evans, 2012). The formation of stars is regulated not only by gravity, but also by magnetic fields (B-fields), turbulence, and radiative and mechanical feedback from nearby stars (McKee & Ostriker, 2007; Hennebelle & Falgarone, 2012; Hennebelle & Inutsuka, 2019). Turbulence counteracts gravity and prevents the collapse of the dense core, but it can also generate shocks that compress the gas and trigger the formation of new stars.
B-fields are thought to play a critical role in the star formation process by supporting the material in the cloud against self-gravitational collapse (see, e.g., Mouschovias, 2001; Crutcher, 2012; Pattle & Fissel, 2019; Pattle et al., 2023). The collapse can occur when the dense core is magnetically super-critical, with the ratio of mass to magnetic flux (known as mass-to-flux ratio) greater than one, i.e., when B-fields are not strong enough to prevent the gravitational collapse. The collapse of super-critical cores will lead to the formation of protostars. Conversely, when the core is sub-critical, strong B-fields can act against gravitational collapse. However, several mechanisms might enhance the mass-to-flux ratio and enable the core to collapse, including ambipolar diffusion (see e.g., Mestel & Spitzer, 1956; Mouschovias et al., 2006; Liu et al., 2022) and magnetic reconnection diffusion (Lazarian & Vishniac, 1999; Santos-Lima et al., 2012).
Feedback from massive stars (O- or B-type) can influence the properties of the parental molecular cloud and the star formation processes (e.g., Krumholz et al., 2014; Pabst et al., 2019, 2020). Three-dimensional (3D) magnetohydrodynamic (MHD) simulations have suggested that B-fields also influence the structure of clouds along with radiation pressure feedback from massive stars (e.g., Henney et al., 2009; Mackey & Lim, 2011a). It has been shown that when the B-field orientations are parallel to the ionizing radiation field, the cloud will tend to have a flattened shape, while the cloud will have a column-like or pillar structure if the B-fields are perpendicular to the radiation field (Mackey & Lim, 2011b). These findings from simulations have been in agreement with observations and could explain various globular and elongated substructures in the IC 1396 region (Soam et al., 2018), the famous pillar structure in the M16 region (Pattle et al., 2018), and the Horsehead Nebula (Hwang et al., 2023).
Interstellar dust grains are known to have aspherical shapes and can be systematically aligned with a fixed direction in space, e.g., interstellar B-fields (Hall, 1949; Hiltner, 1949). The alignment of interstellar dust grains with B-fields involves the internal alignment of the grain axis of the maximum inertia moment (also the shortest axis) with its angular momentum (i.e., internal alignment) and the alignment of the angular momentum with B-fields (external alignment). Internal alignment is caused by internal relaxation processes, including Barnett relaxation and inelastic relaxation (Purcell, 1979). The external alignment is caused by the Radiative Torque (RAT) alignment mechanism (Draine & Weingartner, 1997; Lazarian & Hoang, 2007) and by the more comprehensive magnetically-enhanced RAT mechanism (MRAT, Hoang & Lazarian, 2016). The internal and external alignment processes establish the alignment of the grain with its minor axis parallel to the B-fields (see, e.g., Hoang et al., 2022). Dust grains absorb or scatter starlight in the ultraviolet (UV), optical, or near-infrared (NIR) range. However, aligned dust grains tend to absorb/scatter the electric field component of the starlight along their major axis more efficiently. This results in the polarization of background starlight with the polarization vector parallel to the grain’s minor axis. Thus, the polarization of the starlight seen in the UV to optical or NIR wavelength has the polarization orientation parallel to the B-field orientation. Aligned dust grains also re-emit the energy absorbed from starlight in the longer wavelength from far-infrared (FIR) to sub-millimeter (sub-mm), so-called thermal dust emission. As the electric field component of starlight is absorbed more along the grain’s major axis, the dust grain will re-emit more energy with the electric field component along this major axis than along the minor axis. This causes the thermal emission of aligned dust grains to be polarized with the polarization orientation along the grain’s major axis. Because dust grains have their minor axes aligned with the B-field orientation, the polarization orientation of thermal emission is perpendicular to the B-field orientation. Thus, if the position angle of the polarized thermal dust emission is known, one can rotate the polarization angle by 90° to infer the position angle of the B-fields in the medium. Linear polarized thermal dust emission has been routinely used to trace B-fields in various environments of molecular clouds in the Milky Way (see, e.g., Ward-Thompson et al., 2017; Chuss et al., 2019; Pattle et al., 2022) and extra-galactic regions (see, e.g., Lopez-Rodriguez et al., 2021, 2022; Borlaff et al., 2023). The Davis-Chandrasekhar-Fermi method (DCF; Davis, 1951; Chandrasekhar & Fermi, 1953) has been widely used to measure the strengths of the B-field projected in the plane-of-the-sky, , based on the assumption that the kinetic energy of turbulent gas balances to that of turbulent B-fields.
Adopting the mean value of B-field strengths along the line-of-sight () measured using the Zeeman technique over different molecular clouds, Crutcher et al. (2010) derived a relationship between the B-field strengths and gas volume densities (). For low densities of 300 cm-3, B-fields were found not to considerably change with densities. However, for high densities above 300 cm-3, the B-field strengths increase with densities as , where the power-law index is 0.65. Pattle et al. (2023) compiled a large number of measurements of the B-field strength projected in the plane-of-the-sky () estimated using the DCF method and showed that these generally follow Crutcher et al. (2010)’s results until cm-3. Beyond this value, the distribution is more sparse (see Figure 2a in Pattle et al., 2023). Theoretical studies have suggested that the power-law index varies between 2/3 for weak B-fields (Mestel, 1966) and 0.5 for strong B-fields (Mouschovias & Ciolek, 1999). Therefore, the maps of the morphology and strength of the B-fields within a specific cloud are essential to understanding the role of the local B-fields on star formation within the cloud. Previous authors attempted to derive the map of B-fields toward some well-known star-forming regions, for instance, OMC-1 (Guerra et al. 2021, Hwang et al. 2021), 30 Doradus (30 Dor, Tram et al. 2023), and the G11.11–0.12 filament (Ngoc et al. 2023). These studies revealed significant variations in the B-field morphology and strength across the clouds and demonstrated that B-fields strongly influence the cloud’s evolution and, thus, the star formation activities.
In this work, we focus on a specific system of the Ophiuchus-A (hereafter Oph-A) molecular cloud to study the effects of stellar feedback and B-fields on the evolution of the cloud and star formation. The Oph-A molecular cloud is located in the northern part of the L1688 active low-mass star-forming region in the Ophiuchus dark cloud complex (Loren et al., 1990; Wilking et al., 2008; Esplin & Luhman, 2020). The Oph-A cloud was initially identified as a dense gas clump based on DCO+ observations performed by Loren et al. (1990) among other clumps named from Oph-B to Oph-F. This region has high extinctions (20 mag) corresponding to the high gas column densities compared to other regions, which are typically more quiescent (see Figure 1). Being one of the closest low-mass star-forming regions in our Galaxy, located at a distance of 137 pc (Ortiz-León et al., 2017), the gas and dust content in the Oph–A region has been studied extensively at almost every wavelength by both ground-based and space instruments for many years (see e.g., Wilking et al., 2008, for a review). The Oph-A region hosts numerous prestellar and protostellar cores identified by the Ground Belt Survey (Ladjelate et al., 2020) and protostars in the early stages (i.e., Class 0/I Young Stellar Objects (YSOs), Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010). The Oph-A cloud is mostly influenced by a nearby embedded early high-mass B3/4 star (Brown & Zuckerman, 1975; Mookerjea et al., 2018), Oph-S1, which has a mass of 8 (Hamaguchi et al., 2003). The Oph-S1 star is surrounded by a compact H ii region (Andre et al., 1988) and is believed to trigger the star formation activity in Oph–A, especially along the ridge of the dense cloud with indications of protostellar clumps (see Figure 2a in Motte et al., 1998). The star formation activity in the Oph-A is also believed to be affected by the Sco OB2 association located at the western part of the cloud at a distance of 145 pc (de Zeeuw et al., 1999).
B-fields in Oph-A were investigated using data observed in various wavelengths: the NIR (JHK bands) using SIRPOL on the Infrared Survey Facility (IRSF) 1.4–m telescope (Kwon et al., 2015), FIR by the on the 2.7–m Stratospheric Observatory For Infrared Astronomy (SOFIA)/HAWC (Santos et al., 2019), sub-mm by the 15–m James Clerk Maxwell Telescope (JCMT)/POL–2 (Kwon et al., 2018). Santos et al. (2019) performed the analysis on polarimetric data from SOFIA/HAWC observations at 89 and 154 m to study the change in the slope of the polarization spectra with respect to the column densities and dust temperatures in Oph-A. Later, Tram et al. (2021) used a similar SOFIA/HAWC dataset to investigate the relation between the polarization fraction of thermal dust emission versus dust temperatures in the cloud and to understand the mechanism of dust grain alignment and disruption induced by RATs. Kwon et al. (2018) estimated the values in Oph-A using the JCMT/POL–2 data at 850 m toward a few densest parts of the cloud, yielding values in the range of 0.5–5 mG. The highest strength was measured toward the densest part of the cloud associated with the starless core SM1 (see Figure 1). These strengths in Oph-A are comparable to values found in the OMC region (Hwang et al., 2021). Kwon et al. (2018) also found that the B-field direction is generally perpendicular to the main structure of the cloud, which is generally seen in the dense cores or high-density structures of molecular clouds. The Oph-A region provides an ideal isolated system for studying the impact of the B-fields in the star-forming region and investigating the interplay between B-fields and the stellar feedback into the molecular cloud.
In this paper, we aim to:
1.
determine the B-field morphology and strengths in the entire Oph-A region using the SOFIA/HAWC dust polarization data;
2.
investigate the role of B-fields relative to gravity and turbulence in the star formation activity in the cloud;
3.
quantify the combined effect of B-fields and radiation pressure from the nearby Oph-S1 star on the shape of the Oph-A cloud.
The paper is organized as follows. We describe a set of archival observational data in Section 2. We present the results of mapping the B-field morphology and strength using various techniques in Section 3 and discuss the roles of B-fields relative to gravity, turbulence, and radiation pressure feedback from the Oph-S1 star in Section 4. Finally, we summarize our results in Section 5.
2 Observational Data
2.1 Dust polarization data from SOFIA/HAWC
We use the FIR polarimetric observations toward Oph-A from the High-resolution Airborne Wideband Camera-plus (HAWC; Harper et al., 2018) onboard SOFIA. The observations were done on 2017 May 17, under the proposal ID (PI: Darren Dowell), using the nod-match-chop mode, with 154° chop angle, 480″ chop throw, and 4-points dithers. Here, we use the SOFIA/HAWC Level–4 data for two bands C and D, centered at 89 and 154 m, respectively, which is fully calibrated and can be directly downloaded from the SOFIA archive111https://irsa.ipac.caltech.edu/Missions/sofia.html. The final polarimetric maps cover an area of toward the central region in Oph-A with the angular resolutions of 7.8 and 13.6 full-width-half-maximum (FWHM) at 89 and 154 m, respectively. The pixel size of HAWC maps is 2.0″ at 89 m and 3.4″ at 154 m. The observations and data reduction are described in detail in Santos et al. (2019).
From the maps of the Stokes-, Stokes-, Stokes- and their uncertainties , , , the maps of polarization fraction () and its corresponding uncertainty () are calculated. The de-biased polarization fraction, , is calculated as: (Vaillancourt, 2006). The maps of polarization angle () and its corresponding uncertainty () in degrees can be derived from Stokes- and Stokes-. We refer the reader to Gordon et al. (2018) for a full description of estimates of the polarization fraction, polarization angle, and their corresponding uncertainties. To enhance the quality in measuring the maps of B-field orientation and strength in Oph-A, we apply some critical conditions in choosing vectors (pixels) in SOFIA/HAWC maps. We mask out from SOFIA/HAWC maps the pixels for which ¡54 for band C and ¡214 for band D. In both bands, we further remove the pixels for which ¡3 and ¿50% from HAWC maps (see Appendix A).
2.2 Spectroscopic data
We use the spectroscopic observations of HCO+(43) line at 356.734 GHz obtained from the Heterodyne Array Receiver Program (HARP; Buckle et al., 2009) mounted on JCMT. The observations were done during June and July 2011 (Proposal ID: M11AU13), covering the Oph-A region with a beam size (or FWHM) of 14. We use the reduced spectral map of HCO+(43) line directly downloaded from the JCMT science archive222https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/. The data cube has an original spectral resolution of 0.025 km s-1 and is smoothed to have a final spectral resolution of 0.5 km s-1. For comparison to the HCO+(43) data, we also use observations of N2H+(32) line at 279.515 GHz obtained from the APEX-2 instrument on the 12-m Atacama Pathfinder EXperiment telescope (APEX; Güsten et al., 2006). The data cube has an original spectral resolution of 0.082 km s-1. Here, we use the back-end data cube with a spectral resolution of 0.25 km s-1, covering the Oph-A region with the beam size of 22 (for details, see Liseau et al., 2015).
2.3 Dust temperature and column density maps
We adopt the maps of dust temperature () and gas column density () from Santos et al. (2019). These maps were obtained from the fitting of a modified black body to the spectral energy distribution (SED) using the continuum flux maps from Herschel/PACS (Poglitsch et al., 2010) at 70, 100, and 160 m and have the FWHM of 11.4 (for details, see Section 3.2 in Santos et al., 2019). Figure 1 shows the spatial distribution toward Oph-A, overlaid with contours of column density . In this study, we will use the map of as the map of the gas kinetic temperature, assuming that the gas is in the local thermodynamic equilibrium (LTE) and well-coupled with the dust.
3 Magnetic fields in Oph-A
3.1 Magnetic field morphology
We construct the B-field morphology in Oph-A by rotating polarization angles by 90°. Figure 2 shows the maps of B-field orientation (half-vectors) inferred from maps of the polarization angle at 89 and 154 m . The background color maps show the Stokes- intensity at the respective wavelength. Here, the position angle of B-fields () is measured east of north, and in the range between 0 and 180. The polarimetric data at both wavelengths can well reveal B-fields toward the dense region in Oph-A where Stokes-¿0.12 Jy/arcsec2 (the area inside the lowest contour in each map). The data at 89 m can further resolve the B-fields morphology extended toward the warmer dust region located at the east side of the Oph-S1 star.
The B-field vectors in Oph-A are generally well-ordered and largely follow the material structure in low- to intermediate-densities regimes (see Figure 2). B-field orientations seem to be bent toward the higher-density region in the center of the map and are mostly perpendicular to the ridge of the cloud. On the eastern side of the map, the orientation of B-field vectors tends to vary in the counter-clockwise direction from North to South, i.e., the position angle of B-fields increases. Particularly, in the North-East part of the map, B-field vectors run along the North-East to South-West direction with a median value of 56° and 49° at 89 and 154 m, respectively. This feature is also seen in a small area in the North-West part of the map. Toward the East side, B-field vectors orient nearly to the East-West direction, with median values of 83° at 89 m and 75° at 154 m. On the other hand, B-field vectors in the South-East region tend to orient consistently to the South-East to North-West directions, with median values of 92° and 85° at 89 and 154 m, respectively. Interestingly, the data at 154 m shows B-field vectors running nearly vertically, i.e., along the North-South direction toward a small area at the North part of the map, which is not seen in the data at 89 m (see Figure 2 and also in Figure 15).
In addition, there are two outstanding components of B-fields toward the regions characterized by the highest H2 column densities: one component associated with the peak of continuum emission at 154 m runs along North-East to South-West direction with the median B-field orientation of 70; the other is located at the North of the first component and associated with the peak of continuum emission at 89 m, which exhibits the B-field vectors along the East-West direction with a median position angle of 88° and 81° at 89 and 154 m, respectively (see Figure 15). We will further discuss the B-field morphology together with B-field strengths in Section 4.1.
3.2 Measurement of magnetic field strength using the DCF method
We first use the classical DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953) (hereafter we refer to as the DCF method) to calculate the strengths of the B-field projected in the plane-of-the-sky () in the entire region of the Oph-A cloud. We will compare our obtained results with those obtained using more advanced techniques in Section 3.3.
According to the DCF method, the B-field strength can be estimated by Ostriker et al. (2001):
where is the correction factor, is the gas mass density in units of g cm-3, =2.8 is the mean molecular weight per hydrogen molecule (Kauffmann et al., 2008), is the mass of the hydrogen atom, is the gas volume density in units of cm-3), is the one-dimensional (1D) non-thermal velocity dispersion along the line-of-sight of the ideal gas species that traces the dense region as similar as by dust polarized emission in units of km s-1, and is the polarization angle dispersion in degrees. The correction factor is uncertain, and numerical simulations by Ostriker et al. (2001) show when the polarization angle dispersion does not exceed 25. However, varying values of will result in different strengths (see e.g., Skalidis et al., 2021; Liu et al., 2021, for detail of discussions).
Here, we use the formula as appeared in Crutcher (2004):
where is non-thermal FWHM velocity dispersion measured in units of km s-1, and the correction factor has been used. This formula is widely used to estimate via the DCF technique.
3.2.1 Polarization angle dispersion maps
It is essential to see how the B-field orientation changes locally within the molecular cloud. To do that, we calculate the dispersion in the polarization angle distribution () across the Oph-A region. We use the method presented in Hwang et al. (2021) for which is calculated pixel-by-pixel throughout the map of Oph-A. We summarize the procedure of calculating toward the pixel in the map as follows:
1.
We work with a box covering the pixel with a size of 99 pixels, corresponding to 2 beam sizes of the SOFIA/HAWC observations (i.e., 15.6 and 36.4 in bands C and D, respectively).
2.
We calculate the mean value of polarization angles within the small box, , which implies the local mean polarization angle toward the pixel .
3.
For each pixel within the small box, we calculate the difference between the observed polarization angle and the mean value. We compute the root mean square (RMS) of the distribution of angle differences in the box. This RMS value indicates the polarization angle dispersion calculated toward the pixel .
4.
The uncertainty of the polarization angle dispersion toward the pixel is propagated from the uncertainty of polarization angles in the small box and determined as the mean value of polarization angle uncertainties within the small box.
By repeating the process for all pixels in the entire region, we obtain the map of in Oph-A. We note that if there are more than 50 of pixels undefined (i.e., having NaN value of ) in the computing box, the result of polarization angle dispersion calculation toward pixel is ignored. We apply this criterion to keep the high quality of the estimate. It is also necessary to de-bias the polarization angle dispersion due to the uncertainty of the observed polarized angles.
Figure 3 shows the maps of polarization angle dispersion toward Oph-A using HAWC data in two bands C (left panel) and D (right panel), respectively. The polarization angle dispersion calculated using data in bands C and D are in ranges 2.3–32.2 and 1.3–31.5, which correspond to the median values of 5.7° and 5.8°, respectively. The differences of in two wavelengths imply that the data at the shorter wavelength tends to trace warm dust surrounding the high-mass star S1 and toward the outer layers; whereas the data at the longer wavelength probes the colder dust in the denser layers of the cloud. We note, however, that in maps of in both bands, there are a few pixels where (27 pixels in band C and 45 pixels in band D). Thus, we do not use the value in these pixels to estimate in Section 3.2.4.
3.2.2 Velocity dispersion map
We use the HCO+(43) molecular data observed by JCMT/HARP to estimate the total velocity dispersion, . This molecular line has a critical density of 106 cm-3 (Greve et al., 2009) and thus is considered a good tracer to probe the dense region in the Oph-A cloud. The HCO+(43) emission is well-represented by a single Gaussian profile. Thus, in each pixel, we fit the line with a single Gaussian. The central velocity () across Oph-A covers the range from 2.25 to 3.76 km s-1 with a median value of 3.12 km s-1, which is in agreement with the systemic velocity of Oph-A of 3.44 km s-1 (André et al., 2007). We use the line width of the Gaussian as the measure of total velocity dispersion toward each pixel.
To construct the map of the 1D non-thermal (turbulent) velocity dispersion (), we subtract the thermal component () from the map of the total velocity dispersion, given by:
where =29 amu is the mass of the HCO+ molecule, is the Boltzmann constant, and is the kinetic temperature of the gas. Here, we assume that the gas is in LTE condition and well-coupled with the dust in the cloud. Thus, the gas kinetic temperature is equal to the excitation temperature in all levels () and equal to the dust temperature: . We convolve the map of to the lower resolution of the molecular line map (14) to obtain the same pixel size on both maps. We convert the velocity dispersion of the non-thermal component into the FWHM velocity dispersion, given as . The uncertainty of non-thermal FWHMs is propagated from the uncertainty of the line width obtained from the Gaussian fit.
Figure 4 shows the map of the 1D non-thermal FWHM of the HCO+ molecular line toward Oph-A. The values are in the range from 0.62 to 2.98 km s-1 with a median value of 1.19 km s-1. We find a good agreement between our result obtained with HCO+(43) and those from other line tracers, i.e., 0.95 km s-1 using N2H+(32) from APEX and 1.5 km s-1 using C18O(32) from JCMT/HARP (see Appendix C for details). Thus, it is plausible to use results from the HCO+(43) line in the following analysis.
3.2.3 Gas volume density map
We estimate the gas volume, , across the dense cloud from . Here, we assume that the cloud has a spherical Plummer-like shape (Liseau et al., 2015; Santos et al., 2019), which is frequently used to describe the globular dense cores (Plummer, 1911). We consider a sphere centered at the peak position of , located at (RA, Dec)= (, ), which is also associated with the position of the starless core SM1 (the separation is much less than 1″). We take the radius of the cloud (0.074 pc, Santos et al., 2019). To describe the effect of stellar feedback from the high-mass Oph-S1 star, we further add a spherical shell with a radius of surrounding this star. We also assume that the positions of peak column density and the S1 star are located in the same plane perpendicular to the line of sight. The top panels in Figure 5 show the 3D schematic view of the cloud (left) and a cross-section on the -plane at a given position (right). The - and -axes perpendicular to each other are in the plane of the sky and the axis is along the line of sight. The center of the sphere is located at . The depth of the cloud at each position () along axis is calculated by:
where and are the radius of horizontal circles perpendicular to the plane of sky at position (see top right panel in Figure 5). We obtain the map of , shown in the bottom left panel in Figure 5. The values of are in the range of . We then infer from using the formula:
The bottom right panel in Figure 5 shows the spatial extent of across Oph-A, overlaid by the contours representing the continuum emission in band D (Stokes-). The map has an angular resolution of 11.4″and a pixel size of 3.2, similar to that of the map. The gas density spans from cm-3 and is highest at the center of the cloud. The pattern of distribution well follows the intensity map in band D, revealing the ridge shape of the dense cloud. Some pixels located at the circular boundary in the map show anomalously high densities; this is due to the very low depths caused by the spherical model, and thus we exclude these pixels from our strength measurement.
3.2.4 Magnetic field strength maps
We use maps of polarization angle dispersion (), non-thermal FWHM velocity dispersion (), and gas volume density () to measure in the entire region of Oph-A. All input maps are smoothed to the angular resolution of 14, corresponding to that of the map. We then reconstruct the maps of at 89 and 154 m toward Oph-A using Eq. 2, which are shown in Figure 6.
The values vary from mG and from mG for data at 89 and 154 m, respectively. The highest strength is found at the center region of the cloud, associated with the emission peak at 154 m and also the position of the SM1 core. The median values are and mG333The uncertainties of these median values are solely of the estimated distribution across the cloud. , as measured using 89 and 154 m data, respectively. The uncertainty of is propagated based on the uncertainty of the velocity dispersion of the gas and polarization angle dispersion. Since map is inferred from resulted in SED fitting models (see Santos et al., 2019). In addition, the estimated depth of the cloud is based on the 3D geometrical shape of the cloud and also is uncertain. Thus, we do not take into account the uncertainty of the gas volume density in the total uncertainty of the B-field strengths. For this reason, the uncertainty of might be underestimated. Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : . For the following analysis, we will employ only the map inferred from the polarimetric data at 154 m since it covers the larger area of the higher-density region in Oph-A than that from the data at 89 m.
3.3 Measurement of magnetic field strength using other methods
Several new techniques have been proposed to improve the DCF method for the last decades. For instance, Hildebrand et al. (2009) introduced the angle dispersion function (ADF) method where the dispersion in polarization angles in Eq. 2 is replaced by the square root of the structure function of polarization angles. This provides an improved measurement of B-field strength but also gives us the ratio of the turbulent to large-scale B-field strengths. Alternatively, Skalidis & Tassis (2021) introduced a modification of the DCF method, hereafter ST, assuming that the mean (ordered) and fluctuation components of the total B-field energy, and , respectively, are not independent. More recently, the differential measure approach (DMA; Lazarian et al., 2022) is considered promising to enhance the accuracy of B-field strength measurements at scales smaller than the turbulence injection scale and takes into account the anisotropic properties of MHD turbulence (given by factor ). For the sake of cross-checking the B-field strength measurement in Oph-A using the DCF method, we re-calculate the mean B-field strengths using ADF, ST, and DMA methods at the two subregions (a) and (e) associated with the densest parts of the cloud (these subregions are named as in Kwon et al. (2018), see also in Figure 15).
3.3.1 ADF method
According to the ADF method, the can be estimated using Eq. 2 but the polarization angle dispersion is replaced by the structure function of the polarization angles, so-called angular structure function (Hildebrand et al., 2009). The angular structure function, , at an angular scale is estimated as:
where denotes an average, is the polarization angle difference between individual pairs of polarization vectors at position and , where and are the corresponding polarization angles at position and , respectively, and is number of pairs of vectors with a displacement of . At scales much smaller than the scale for a variation of the large-scale B-fields structure (typically of 5 seen by Planck observations), the total angular dispersion function can be expressed as:
where represents the turbulent component of B-fields (), describes the large-scale structure of B-fields (), and is subject to measurement uncertainties. Also, the ratio between the turbulent and large-scale B-field strength indicates the angular dispersion, given as (Hildebrand et al., 2009):
When , i.e., rad, one can obtain .
For each subregion (a) and (e) in Oph-A, we calculate the angular structure function using HAWC data at 154 m via Eq. 6 and then perform the fit to the quadratic function described in Eq. 7 using the Python curve-fit routine. We select data points that 13.6″50″ to avoid fitting to structures higher than the beam size and to better fit to smaller-scale structures. Top panels in Figure 7 show structure functions of polarization angles as a function of the distance of displacement toward subregions (a) and (e). From the fitted value of the fits, we derive values of and toward subregions (a) and (e), respectively. We find these values in good agreement with mean polarization angle dispersion values estimated using our approach described in Section 3.2.1, which are 9.2° and 14.3° in subregions (a) and (e), respectively. Toward subregions (a) and (e), mean gas volume densities are and cm-3 respectively, and their mean non-thermal FWHMs are 1.74 and 1.45 km s-1, respectively. We substitute new values together with the above mean values of and into Eq. 2 to re-calculate strengths in subregions (a) and (e), which are mG and mG, respectively. These values are in good agreement with strengths obtained using the DCF method, where values are in ranges of 0.32–1.55 mG and 0.15–0.97 mG toward subregions (a) and (e), respectively, with corresponding mean values of 0.84 and 0.37 mG.
3.3.2 ST method
The ST method is applicable for the compressible B-fields mode and thus is expected to provide a more accurate estimate of B-field strengths (Skalidis et al., 2021). Another advantage of the ST method is that this method accounts for the correlation between the fluctuation and mean components of the total B-field energy not to be zeros. Here, we measure the maps of strength () in band C and D using the formula as appeared in Skalidis & Tassis (2021):
We show the maps of across Oph-A in band C and D in Figure 16. We find that value ranges from mG at 89 m and mG at 154 m, appeared to be lower than . A median value of mG is seen in both bands. For a direct comparison with those using other methods, we estimate the mean value of at 154 m toward two subregions (a) and (e), which are 0.19 and 0.07 mG, respectively. These values are about 4.5 times lower than those estimated using the DCF method.
3.3.3 DMA method
We also re-calculate trengths in two subregions (a) and (e) using the DMA method, , using Eq. 10 following Lazarian et al. (2022):
where and are the second order structure function of polarization angles and central velocity at an angular scale , is the correction factor depending on the anisotropic properties of MHD turbulence. Here, we adopt =1 related to the naive formalism as suggested in Lazarian et al. (2022).
We calculate the structure function of the central velocity of HCO+(43) line toward subregions (a) and (e), using the similar form in Eq. 6 but for the central velocity , shown in middle panels in Figure 7. We then measure strengths in two subregions (a) and (e) via Eq. 10, using the information from the together with and the mean values of in the subregions. The as a function of displacement is shown in bottom panels of Figure 7 for subregions (a) and (e). It can be seen that the resulting in the range is likely flatten within their uncertainty. Thus, in each subregion we take the mean of in this range to be the average value, which are 0.34 mG and 0.12 mG in subregions (a) and (e), respectively. These values are slightly higher than the resulting estimated using the ST method, but at the low-end of that measured using the DCF method, i.e., 2–3 times less than the mean values estimated using the DCF method presented in Section 3.2. These differences are indeed expected since the DCF method somewhat overestimates the strength in the cloud. We will discuss this issue in Section 4.1.
3.4 Mass-to-flux ratio
The relative importance of B-fields compared to gravity in the cloud can be quantified by the mass-to-magnetic-flux ratio (or so-called mass-to-flux ratio), , which can be estimated in units of the critical ratio, (Nakano & Nakamura, 1978): . We measure in the entire region of Oph-A, using the formula introduced by Crutcher (2004):
where and are in units of cm-2 and G, respectively.
Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : . For the following analysis, we will employ only the map estimated using the DCF method and inferred from the polarimetric data in band D since it covers the larger area of the higher-density region in Oph-A than that from the data in band C.
The top panel in Figure 8 shows the spatial distribution of the mass-to-flux ratio in Oph-A. It reveals an increasing gradient of from the central and eastern parts of the clump to the outer regions, spanning from 0.03–2.56, with a median value of 0.26. The region exhibiting the lowest values of 1 is close to the OphS1 star, suggesting that the region is firmly magnetically sub-critical, i.e., B-fields are strong enough to prevent the gravitational collapse. This can also be explained by the influence of the strong winds and radiation fields from the OphS1 star which tend to push the material away to to its vicinity at the west side. The mass-to-flux ratio increases toward the outskirt regions of the Oph-S1 star, especially toward regions along the ridge of the dense cloud (i.e. center, north, and south-east parts of the map) where the gas column density is cm-2. These regions are magnetically super-critical and so, the self-gravitational collapse might have occurred there, resulting in several protostars and a single protostellar core identified in the region (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
3.5 Alfvénic Mach number
To investigate the interplay between B-field and turbulence due to the non-thermal motion of the gas in Oph-A, we calculate the map of Alfvénic Mach number, given as:
where is given by Eq. 3, and is the Alfvénic speed driven by the B-fields (), both in units of km s-1. The factor in Eq. 12 assumes that turbulence is isotropic, and allows the calculation of the three-dimensional (3D) Alfvénic Mach numbers. We note that adding this factor will provide upper limits to the estimated .
The middle panel in Figure 8 shows the map of toward Oph-A using results of polarimetric data at 154 m and HCO+(4–3) as the gas tracer. The values are in the range from 0.1–1.5 across the cloud, with a median value of 0.38. Most of Oph-A shows ¡1 (magnetically sub-Alfvénic), suggesting the B-field dominates turbulence in the cloud. However, two small regions at the north and southwest of the map exhibit 1, indicating strong Alfvénic motions of the gas (magnetically supper-Alfvénic). These regions are also associated with the positions of a few prestellar and protostellar cores, and YSO candidates (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
3.6 Plasma parameter
To compare the magnetic presure () and thermal pressure () in Oph-A, we construct the map of the plasma beta parameter (), given as:
where is the thermal sound speed measured in units of km s-1, assuming = (see Section 3.2.2).
The bottom panel in Figure 8 shows the map of the parameter toward Oph-A in band D. We obtain ¡1 across the map with a median value of 0.02, implying that the cloud is supported by strong magnetic pressure which is much more prominent than the one induced by the thermal motions of the gas.
3.7 Virial mass analysis
To probe the dynamical state of the Oph-A cloud, we calculate the virial parameter (), which is the ratio of the virial mass () and the isothermal mass () of the cloud. Here, we assume the cloud as a uniform density sphere with an effective radius of as a whole, having a mean mass density of . The isothermal mass, , considering the self-gravity of the cloud can be estimated as:
The virial mass, , determined from magnetic, thermal, and non-thermal kinetic energies can be estimated as (see e.g., Bertoldi & McKee, 1992; Pillai et al., 2011):
where , , and are respectively the mean values of the Alfvénic velocity, thermal sound speed, and of non-thermal velocity dispersion toward the cloud, is the gravitational constant.
We calculate by selecting an ellipse covering the area where is measured. This ellipse is centered at (RA, Dec) = () with the semi-major and semi-minor axes are =120 and =75, respectively. The is then determined as (0.063 pc assuming the distance to Oph-A of 137 pc). Within this area, the mean gas volume density is cm-3, corresponding to the mean mass density . The corresponding mean values of velocities are: , , and . We use these mean values of the velocities, non-thermal velocity dispersion, and the gas volume density to calculate and in Eqs. 14 and 15, yielding 15. This shows that the cloud is gravitationally unbound and that B-fields and turbulence can support the cloud against self-gravity. This high value of obtained in the cloud overalls could explain the low efficiency of star formation in rho Oph-A though in a few regions the density is very high up to 105 or 106 cm-3. This is consistent with the lack of prestellar cores and protostellar sources in the central region, which is seen in the lower-density parts of the cloud only (see e.g., Figure 8). In order to study smaller scales across the clump, polarimetric observations with a higher spatial resolution would be required.
4 Discussion
4.1 Morphology and strengths of magnetic field in Oph-A
The B-field orientation in Oph-A is well-ordered and predominantly perpendicular to the ridge of the dense cloud (Section 3.1). The feature is illustrated most clearly by the data at the longer wavelength (154 m). The data at 89 m better traces warmer dust in the lower-density region of Oph-A, which surrounds the Oph-S1 star.
The B-field properties measured using SOFIA/HAWC polarization data are qualitatively consistent with previous measurements of B-field orientations in Oph-A obtained at 850 m using SCUBA-2/POL–2 (Kwon et al., 2018) and SCUBA/SCUPOL (Matthews et al., 2009) on JCMT and at NIR wavelengths (JHK bands) using IRSF/SIRPOL (Kwon et al., 2015). In the POL–2 dataset, Oph-A was divided into 10 subregions, identified based on their B-field position angle and the polarization fraction, covering the entire extent of HAWC maps (see Figure 15). The highest density region is characterized by an average B-field position angle of 50°, which is also consistent with the NIR data toward the L1688 main cloud (Kwon et al., 2015). Table 1 shows the median B-field position angle () in each subregion which has been covered also by the FIR and sub-mm observations. The median B-field position angles from the 89, 154, and 850 m datasets are in rather good agreement, with the difference in values ranging from 3 to 19 (30%). These differences could be explained by the fact that the sub-mm data probe colder dust layers of the cloud, while the FIR data trace warmer layers. Indeed, such foreground layers were identified through the detection of the self-absorption features in the spectra of the [O i] line at 63 m (Liseau et al., 2006) and the [C ii] line at 158 m (Mookerjea et al., 2018).
Subregiona | Diff.b | Diff.b | ||||||
(deg) | (%) | (%) | (mG) | |||||
(a) | 36 | 30 | 4.5 | 2.1 | 5.0 | |||
(b) | 42 | 23 | 0.1 | 0.2 | – | |||
(e) | 12 | 19 | 0.7 | 0.3 | 0.8 | |||
(f) | 21 | 28 | 0.2 | 0.3 | – | |||
(g) | 25 | 14 | 0.5 | 0.7 | – | |||
(i) | 23 | 13 | 0.4 | 0.5 | – |
Notes: (a) The subregions are named as in Kwon et al. (2018). (b) The relative difference in is taken as , where is the B-field position angle at wavelength =[89, 154] (m).
Kwon et al. (2018) calculated in three subregions associated with the densest parts of Oph-A using non-thermal velocity dispersion from the N2H+(1–0) line (André et al., 2007) and the sub-mm data at 850 m. In our study, we calculate pixel-by-pixel toward the entire Oph-A region. To facilitate comparisons with results obtained from the sub-mm data, we calculate the mean value of in two out of these three subregions, covered by the HAWC data (see Table 1). Mean values of using FIR data are lower than that obtained with the sub-mm data, which traces colder dust layers in the cloud characterized also by the higher densities. The HAWC data, on the other hand, essentially traces the warmer layers in the lower-density region. Also, Kwon et al. (2018) adopted values of the polarization angle dispersion as the standard deviation of the polarization angles, which are much smaller than those using the spread of the dust polarization angles (see Section 3.2.1). As a consequence, the obtained with sub-mm data is expected to be stronger than the one obtained with FIR data. We find that our estimate for Oph-A is typically higher than those in other molecular clouds, ranging from 50 to 400 G (for e.g., Coudé et al., 2019; Ngoc et al., 2021; Pattle et al., 2021; Ward-Thompson et al., 2023), but similar to those found in the DR21 (OH) region (Poidevin et al., 2013) or the OMC cloud in Orion A region (Pattle et al., 2017; Hwang et al., 2021) where of a few mG was measured.
In this study, we adopt component as the total B-field strength. If we consider the contribution of the LOS component, the total B-field strength will be stronger than what we measured. The mass-to-flux ratio, Mach Alfvénic number, and plasma parameter will decrease, but this will not significantly change our conclusion that B-fields are generally dominant over gravity, turbulent, and gas kinetic energies, especially toward the center of the cloud. Furthermore, as presented in Section 3.3, we test that the strengths obtained with the DCF method in subregions (a) and (e) are rather in good agreement with values obtained with the ADF method but higher than those using the ST and DMA methods (see Table 2). We use new values estimated with ADF, ST, and DMA methods toward two subregions (a) and (e) to estimate the mass-to-flux ratio, Mach Alfvénic number, and plasma, which are all higher than that measured with the DCF method (see Table 2). Subregion (a) changes from magnetically sub/trans-critical to super-critical but remains to be super-Alfvénic, and has ¡1 in overall. A similar behavior is seen in subregion (e) except that this subregion changes from trans- to super-Alfvénic (see Table 2).
We now assume that the B-field strength in the entire cloud decreases by a factor of 4.5 as measured with the ST method. In this case, the majority of the cloud will become magnetically trans/super-critical (1), except for the regions in the vicinity of the Oph–S1 star (see the top panel of Figure 20). However, this trend of remains similar as obtained using . The cloud will become mostly super-Afvénic with ¿1, except for the southeast region of the cloud (see the middle panel of Figure 20). This assumption will also result in the Alfvénic velocity decreasing by a factor of 4.5 and thus, . This implies that the cloud is in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales.
Subregion | DCF | ADF | ST | DMA | DCF | ADF | ST | DMA | DCF | ADF | ST | DMA | DCF | ADF | ST | DMA |
a | 0.84 | 0.69 | 0.19 | 0.34 | 0.91 | 1.12 | 4.03 | 2.49 | 1.57 | 1.92 | 6.93 | 3.88 | 0.02 | 0.03 | 0.37 | 0.12 |
e | 0.37 | 0.26 | 0.07 | 0.12 | 0.84 | 1.22 | 4.20 | 1.99 | 0.62 | 0.90 | 3.09 | 1.96 | ¡0.01 | ¡0.01 | 0.11 | 0.04 |
4.2 Magnetic field orientation versus density structure of Oph-A
To investigate what is the relative orientation of the B-fields with respect to the gas structure in Oph-A, we quantify the projected offset angle between the orientations of B-field vectors projected in the plane-of-the-sky and the main structure associated with the ridge of the cloud, given as , where is the position angle of the ridge. We identify the shape of the ridge using the Python package RadFil444https://github.com/catherinezucker/radfil (Zucker & Chen, 2018, see appendix F).
Figure 9 shows the map of offset angles () in Oph-A, overlaid by contours of continuum emission at 154 m. Clearly, the offset angle appears to increase when moving from the lower-density regions to the denser regions, indicating that the relative orientation of B-fields with respect to the main structure of the cloud turns from parallel to perpendicular toward the high-density region. This is in agreement with Planck polarization observations at 353 GHz (850 m) toward the entire L1688 region where the relative orientation of B-fields with respect to the gas structure changes gradually from parallel in the low-density region to perpendicular toward the higher-density region, with the transition occurring at cm-2 corresponding to the visual extinction of 3.3 mag555We convert from using the relation: molecules cm-2 mag-1 with a “standard” assumption of the total-to-selective extinction ratio, , of 3.1 (Frerking et al., 1982). (Soler, 2019). This transition is also observed in other nearby molecular clouds in the Gould Belt using Planck 353 GHz polarization observations, typically occurring at cm-2 or 2.7 mag (see e.g., Planck Collaboration et al., 2016; Soler et al., 2017). This phenomenon can be explained by the motion of the high-density material during the self-gravitational collapse. When the contraction occurs, the material becomes denser and compressed, which could cause the B-field lines to be re-configured. Eventually, the B-field lines are dragged and reshaped to become more uniform and perpendicular to the high-density structure of the field.
The HAWC data traces B-fields in the dense region of Oph-A with cm-2 (see Figure 1), which corresponds to in range of 3.2–256 mag. Thus, we further investigate in detail the relation between the relative orientation of B-fields with respect to the cloud structure and the within Oph-A, as shown in Figure 10. Clearly, in the regime where 30 mag the offset angle () decreases with and drops from 60° at =4 mag down to 0° at 20 mag. This indicates that the relative orientation of B-fields changes from perpendicular back to parallel at 18–33 mag. Similar trends of this transition were found in the dense region of the Serpens South filament (Pillai et al., 2020) and in Serpens Main dense cloud (Kwon et al., 2022), occurring at 21 mag and 50 mag, respectively. This is likely because the B-fields are already dragged into the flow of very dense material during the gravitational contraction. Interestingly, we observe that the offset angle then progressively increases with 40 mag in our region, indicating the relative orientation of B-field with respect to the ridge changes back from parallel to perpendicular (with the transition at 42–60 mag). We note, however, that this feature was not seen in the Serpens South filamentary cloud whose maximum 60 mag (see Figure 3 in Pillai et al. (2020)) but in the Serpens Main cloud region where 170 mag (see Figure 6 in Kwon et al. (2022)).
4.3 Magnetic field versus gas density
The B-field strength ( or ) can be expressed as a function of (): . Theoretical models predict 2/3 in the weak B-field regions (Mestel, 1966) and 0.5 in the strong B-field regions where magnetic energy dominates over the gravity (Mouschovias & Ciolek, 1999). Consequently, in the latter case, it suggests that would increase faster than if there is any collapse in the dense cloud.
Here, we show in Figure 11 the relation toward the Oph-A cloud, using B-fields measured with the DCF method in band D. We find that for densities , the B-fields distribution tends to be flattened. In a regime with densities higher than we start seeing the increasing trend of B-fields with respect to gas densities. Thus, we decide to test the relation at the densest part of the cloud only. We choose this region as the area within the third contour of Stokes- at band D (see in the bottom right panel of Figure 5), where Stokes. Within this region, it reveals a relatively weak correlation between the B-fields and densities (with the Pearson coefficient of the correlation is found , corresponding to a significance of ). We perform the power-law fit using the Python curve-fit routine (), using mean values of estimated within every interval of =0.08666The bin width is chosen by using the Freedman-Diaconis rule. , which results in a power-law index of . In Figure 11, we also overplot the results of using other methods in subregions (a) and (e). We find that the fitted line well covers the data point representing the resulting value using ADF method and shows good agreement with those estimated with ST and DMA in subregion (a) within of uncertainty. The resulting is found consistent with theoretical predictions for strong B-fields (Mouschovias & Ciolek, 1999). This is also consistent with , representing strong B-fields in Oph-A (see bottom panel in Figure 8 and Section 3.6). We further perform the fit to values measured with the ST method from Sect. 3.3.2 and find . We find a similar power-law index for the dataset at band C, which are and using result of and , respectively (see Appendix E).
4.4 Magnetic fields versus stellar feedback
High-mass stars irradiate, heat up, and ionize the surrounding gas and create the H ii region (see e.g., Zinnecker & Yorke, 2007). The ionized, hot gas also expands and interacts with the vicinity environment. The radiation pressure acting on both dust and gas (assuming dust is well-coupled to the gas in the cloud) can drive the material away into the denser region. The lower-density matter in the outer regions of Oph-A is more exposed to the ionization front and pushed away faster than the higher-density matter in the inner regions. This results in dust and gas being compressed and forming the dense shell structure at the edge of the cloud. Here, we investigate a possible impact of the radiation field from the high-mass star Oph-S1 on the Oph-A cloud. The Oph-A molecular cloud is directly exposed to the strong UV radiation field emitted by the Oph-S1 star from the East. Figure 12 shows the map of the continuum emission at 154 m from SOFIA/HAWC+, superimposed by the white contours showing the H2 column density.
We calculate the external radiation pressure of the Oph-S1 star acting on dust and gas in the densest part of the cloud, given by:
where =1100 is the luminosity of the Oph-S1 star adopted from Lada & Wilking (1984) and 0.06 pc is the projected distance to the Oph-S1 star. We obtain K cm-3, which is in agreement with the result calculated from the ionizing photons emitted by the Oph-S1 star (2 K cm-3, Pattle et al., 2015). Here, we omit the fact that the UV radiation from the Oph-S1 star can be absorbed by intervening dust in between the star and the Oph-A region. Thus, the actual acting on dust and gas in Oph-A is overestimated in the Eq. 16. To compare with the external radiation pressure, we calculate the internal thermal pressure due to the hot gas in Oph-A, . We consider the hot gas within the densest part of Oph-A where ¿0.4 Jy/arcsec2 at 154 m only, which corresponds to the area inside the third contour shown in bottom right panel of Figure 5. Within this region, we adopt in the range of 104–106 cm-3 and the average temperature of the hot gas 20 K, which yields in the range of (0.2–20) K cm-3. With this assumption, we note that could be underestimated since there is a gas heating process in the cloud dominated by the UV radiation from the Oph-S1 star and by the photoelectric emission from dust exposed to the UV radiation (i.e., ). The highest internal thermal pressure of 20106 K cm-3 is seen toward the densest part of the cloud, associated with the position of the SM1 core. This suggests that the internal thermal pressure is strong enough to help the material refrain from external pressure. We note that, however, in the vicinity region away from the center (for e.g., the tails of the ridge shown in Figure 12), the internal thermal pressure is weaker, at least one order of magnitude less than the external one. In this case, the radiation pressure from the Oph-S1 star on the east side is dominant and able to push away all the matter into the surrounding environment toward the west side. However, it is clear that the material in the low-density region still survives, despite owning a low internal thermal pressure (see Figure 12), indicating that there must be another mechanism at play, for example, the presence of strong B-fields of a few mG in the cloud. We estimate roughly the magnetic pressure () to be 107–108 K cm-3 for B-field strengths between 1–2 mG. This indicates that the magnetic pressure is at least two orders of magnitude higher than the external pressure of 106 K cm-3, sufficient to support the matter against the external pressure along with the gas thermal pressure.
We further investigate the interplay between B-fields and radiation directions on the evolution of the cloud. We determine the main radiation direction from the Oph-S1 star toward the densest region associated with the position of the starless core SM1 as the sky-projected direction from the Oph-S1 star and SM1 core. We find that the main radiation direction is 70°, measured east of north. The main B-field orientation toward the densest region of Oph-A of 74°, almost parallel to the radiation field. The strong radiation pressure from the Oph-S1 star pushes all the material away and creates a dense shell toward the west side. To prevent dust and gas compression in the dense region, the B-fields exert a magnetic pressure perpendicular to the B-field orientation, seen perpendicular to the direction of the radiation pressure. This suggests that the magnetic pressure does not help the matter resist the external pressure, but tends to stretch the structure out along the northwest-southeast direction, resulting in an elongated shape of the dense shell. In the west side, the external thermal pressure is blocked by the gas thermal pressure due to the high-density material, resulting in the ridge shape in the cloud with very high densities. These results are consistent with results from 3D MHD simulation, which predicted that if the B-field orientation is parallel or along with the radiation direction, the cloud will tend to have a flattened shape Henney et al. (2009). However, in the northeast and southeast regions the offset angles between the B-field vectors and radiation field are larger, i.e., the B-field orientation becomes perpendicular to the radiation direction. In this case, the magnetic pressure and tension induced by B-fields run parallel with the radiation pressure, suggesting magnetic pressure can support the matter to confront the radiative compression and prevent the sweeping dust and gas away. This results in bending the material structure and forming a dense ridge with an umbrella-like shape. Interestingly, in the case study of Oph-A the ionization source is very close to the cloud (¡0.1 pc) and the offset position angle between the radiation and B-fields is not homogeneous but varies from parallel to orthogonal. This allows us to examine widely how the interplay between the B-fields and radiation fields could influence the material structure of the cloud.
5 Summary
In this paper, we presented the analysis using the SOFIA/HAWC polarimetric observations toward the Oph-A molecular cloud in the L1688 active star-forming region of the Ophiuchus dark cloud complex. Our main results are as follows:
Using the dust polarization maps at 89 and 154 m, we derived the map of B-field morphology in Oph-A. We found that B-fields are generally well-ordered and primarily perpendicular to the ridge of the cloud toward the higher-density regime. We estimated in the entire region of Oph-A using the DCF method, ranging from 0.2 to 2.3 mG at 89 and to 2.5 mG at 154 m. We found is prominent toward the densest region of the cloud associated with the starless core SM1 and decreases toward the outskirts. This is the first time we estimate the distribution of the B-field strengths in Oph-A, which is essential to understanding how B-fields vary in the molecular cloud globally. Our derived B-fields morphology and strengths well agree with previous measurements using the sub-mm dust polarization (Kwon et al., 2018). We also estimated strength using more advanced techniques, including the ADF, ST, and DMA, toward two subregions associated with the densest part of the cloud. While strengths estimated with the ADF method were in agreement with those from DCF, the ST and DMA methods provided lower values of 4.5 and 2-3 times, respectively.
We studied the role of B-fields relative to gravity and turbulence by calculating the maps of the mass-to-flux ratio, Alfvénic Mach number, and plasma parameter toward Oph-A using the inferred from SOFIA/HAWC data at 154 m. We found that toward the central dense parts the cloud is magnetically sub-critical and sub-Alfvénic where B-fields dominate over the gravity and turbulence. We obtained in Oph-A, suggesting the cloud is supported by the strong magnetic pressure and much greater than the thermal gas pressure. The lower strengths obtained with the ST and DMA methods do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation. Our virial analysis in the entire cloud showed the virial parameter 15, suggesting that the cloud is gravitationally unbound and able to against the self-gravitational collapse. The cloud may be in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales. We should notice that our rough estimation of is a global parameter, applied for the entire cloud. In contrast, Oph-A is known as one of the densest parts of the active star-forming region L1688 hosting a few pre- and proto-stellar cores and embedded protostars. Higher spatial resolution observations toward the smaller scale region in the cloud will be necessary to investigate in detail the dynamical state of the cloud.
We examined the role of B-fields in supporting the cloud against the radiation pressure from the nearby high-mass star Oph-S1. We found that at the center part of the cloud, the magnetic pressure indeed does not play a role in helping the material prevent radiation pressure. However, toward the northeast and southeast regions of the cloud, the magnetic pressure positively supports the cloud to prevent the radiative pressure together with the gas thermal pressure.
Acknowledgements.
We thank our anonymous referee for their careful reading of our manuscript and valuable suggestions, which improved the quality of this work, particularly with regard to the realistic estimate of the gas volume density in the Oph-A cloud. We thank Bhaswati Mookerjea for the valuable comments and discussion at the early stage of the project and for kindly sharing the final data cube of the HCO+(4–3) transition line. We thank Fabio Santos for sharing the maps of dust temperatures and gas column density. We also thank Alex Lazarian for the constructive discussions concerning the DCF and DMA methods. N.L. thanks Thang Nguyen Tat for his valuable comments as part of the 3D structure of the Oph-A. N.L. and A.K. acknowledge support from the First TEAM grant of the Foundation for Polish Science No. POIR.04.04.00-00-5D21/18-00 (PI: A. Karska). A.K. acknowledges also support from the Polish National Agency for Academic Exchange grant No. BPN/BEK/2021/1/00319/DEC/1. N.B.N. was funded by the Master, Ph.D. Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2022.TS083. M.H. acknowledges support by the National Science Centre through the OPUS grant no. 2015/19/B/ST9/02959. N.L. also acknowledges support by the grant from the Simons Foundation to IFIRSE, ICISE (916424, PI: N.H.). P.N.D., N.B.N., and N.T.P. were funded by the Vietnam Academy of Science and Technology under project code THTETN.03/24-25. This work is based on observations made with the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA). SOFIA is jointly operated by the Universities Space Research Association, Inc. (USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50 OK 2002 to the University of Stuttgart. This work made use of data from the JCMT/HARP and APEX. JCMT is operated by the East Asian Observatory on behalf of National Astronomical Observatory of Japan; the UK STFC under the auspices of grant number ST/N005856/1; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the Operation, Maintenance and Upgrading Fund for Astronomical Telescopes and Facility Instruments, budgeted from the Ministry of Finance of China. This work used the facilities of the Canadian Astronomy Data Centre (CADC) operated by the National Research Council of Canada.References
- André et al. (2007) André, P., Belloche, A., Motte, F., & Peretto, N. 2007, A&A, 472, 519
- Andre et al. (1988) Andre, P., Montmerle, T., Feigelson, E. D., Stine, P. C., & Klein, K. L. 1988, in Activity in Cool Star Envelopes, ed. O. Havnes, B. R. Pettersen, J. H. M. M. Schmitt, & J. E. Solheim, Vol. 143, 293
- Bertoldi & McKee (1992) Bertoldi, F. & McKee, C. F. 1992, ApJ, 395, 140
- Borlaff et al. (2023) Borlaff, A. S., Lopez-Rodriguez, E., Beck, R., et al. 2023, ApJ, 952, 4
- Brown & Zuckerman (1975) Brown, R. L. & Zuckerman, B. 1975, ApJ, 202, L125
- Buckle et al. (2009) Buckle, J. V., Hills, R. E., Smith, H., et al. 2009, MNRAS, 399, 1026
- Chandrasekhar & Fermi (1953) Chandrasekhar, S. & Fermi, E. 1953, ApJ, 118, 113
- Chen et al. (2019) Chen, H. H.-H., Pineda, J. E., Goodman, A. A., et al. 2019, ApJ, 877, 93
- Chuss et al. (2019) Chuss, D. T., Andersson, B. G., Bally, J., et al. 2019, ApJ, 872, 187
- Connelley & Greene (2010) Connelley, M. S. & Greene, T. P. 2010, AJ, 140, 1214
- Coudé et al. (2019) Coudé, S., Bastien, P., Houde, M., et al. 2019, ApJ, 877, 88
- Crutcher (2004) Crutcher, R. M. 2004, in The Magnetized Interstellar Medium, ed. B. Uyaniker, W. Reich, & R. Wielebinski, 123–132
- Crutcher (2012) Crutcher, R. M. 2012, ARA&A, 50, 29
- Crutcher et al. (2010) Crutcher, R. M., Wandelt, B., Heiles, C., Falgarone, E., & Troland, T. H. 2010, ApJ, 725, 466
- Davis (1951) Davis, L. 1951, Physical Review, 81, 890
- de Zeeuw et al. (1999) de Zeeuw, P. T., Hoogerwerf, R., de Bruijne, J. H. J., Brown, A. G. A., & Blaauw, A. 1999, AJ, 117, 354
- Draine & Weingartner (1997) Draine, B. T. & Weingartner, J. C. 1997, ApJ, 480, 633
- Enoch et al. (2009) Enoch, M. L., Evans, Neal J., I., Sargent, A. I., & Glenn, J. 2009, ApJ, 692, 973
- Esplin & Luhman (2020) Esplin, T. L. & Luhman, K. L. 2020, AJ, 159, 282
- Evans et al. (2009) Evans, Neal J., I., Dunham, M. M., Jørgensen, J. K., et al. 2009, ApJS, 181, 321
- Frerking et al. (1982) Frerking, M. A., Langer, W. D., & Wilson, R. W. 1982, ApJ, 262, 590
- Gordon et al. (2018) Gordon, M. S., Lopez-Rodriguez, E., Andersson, B. G., et al. 2018, arXiv e-prints, arXiv:1811.03100
- Greve et al. (2009) Greve, T. R., Papadopoulos, P. P., Gao, Y., & Radford, S. J. E. 2009, ApJ, 692, 1432
- Guerra et al. (2021) Guerra, J. A., Chuss, D. T., Dowell, C. D., et al. 2021, ApJ, 908, 98
- Güsten et al. (2006) Güsten, R., Nyman, L. Å., Schilke, P., et al. 2006, A&A, 454, L13
- Hall (1949) Hall, J. S. 1949, Science, 109, 166
- Hamaguchi et al. (2003) Hamaguchi, K., Corcoran, M. F., & Imanishi, K. 2003, PASJ, 55, 981
- Harper et al. (2018) Harper, D. A., Runyan, M. C., Dowell, C. D., et al. 2018, Journal of Astronomical Instrumentation, 7, 1840008
- Hennebelle & Falgarone (2012) Hennebelle, P. & Falgarone, E. 2012, A&A Rev., 20, 55
- Hennebelle & Inutsuka (2019) Hennebelle, P. & Inutsuka, S.-i. 2019, Frontiers in Astronomy and Space Sciences, 6
- Henney et al. (2009) Henney, W. J., Arthur, S. J., de Colle, F., & Mellema, G. 2009, MNRAS, 398, 157
- Hildebrand et al. (2009) Hildebrand, R. H., Kirby, L., Dotson, J. L., Houde, M., & Vaillancourt, J. E. 2009, ApJ, 696, 567
- Hiltner (1949) Hiltner, W. A. 1949, Nature, 163, 283
- Hoang & Lazarian (2016) Hoang, T. & Lazarian, A. 2016, ApJ, 831, 159
- Hoang et al. (2022) Hoang, T., Tram, L. N., Minh Phan, V. H., et al. 2022, AJ, 164, 248
- Hwang et al. (2021) Hwang, J., Kim, J., Pattle, K., et al. 2021, ApJ, 913, 85
- Hwang et al. (2023) Hwang, J., Pattle, K., Parsons, H., Go, M., & Kim, J. 2023, AJ, 165, 198
- Kauffmann et al. (2008) Kauffmann, J., Bertoldi, F., Bourke, T. L., Evans, N. J., I., & Lee, C. W. 2008, A&A, 487, 993
- Kennicutt & Evans (2012) Kennicutt, R. C. & Evans, N. J. 2012, ARA&A, 50, 531
- Koch & Rosolowsky (2015) Koch, E. W. & Rosolowsky, E. W. 2015, MNRAS, 452, 3435
- Krumholz et al. (2014) Krumholz, M. R., Bate, M. R., Arce, H. G., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 243–266
- Kwon et al. (2018) Kwon, J., Doi, Y., Tamura, M., et al. 2018, ApJ, 859, 4
- Kwon et al. (2015) Kwon, J., Tamura, M., Hough, J. H., et al. 2015, ApJS, 220, 17
- Kwon et al. (2022) Kwon, W., Pattle, K., Sadavoy, S., et al. 2022, ApJ, 926, 163
- Lada & Lada (2003) Lada, C. J. & Lada, E. A. 2003, ARA&A, 41, 57
- Lada & Wilking (1984) Lada, C. J. & Wilking, B. A. 1984, ApJ, 287, 610
- Ladjelate et al. (2020) Ladjelate, B., André, P., Könyves, V., et al. 2020, A&A, 638, A74
- Lazarian & Hoang (2007) Lazarian, A. & Hoang, T. 2007, MNRAS, 378, 910
- Lazarian & Vishniac (1999) Lazarian, A. & Vishniac, E. T. 1999, ApJ, 517, 700
- Lazarian et al. (2022) Lazarian, A., Yuen, K. H., & Pogosyan, D. 2022, ApJ, 935, 77
- Li (2021) Li, H.-B. 2021, Galaxies, 9, 41
- Liseau et al. (2006) Liseau, R., Justtanont, K., & Tielens, A. G. G. M. 2006, A&A, 446, 561
- Liseau et al. (2015) Liseau, R., Larsson, B., Lunttila, T., et al. 2015, A&A, 578, A131
- Liu et al. (2021) Liu, J., Zhang, Q., Commerçon, B., et al. 2021, ApJ, 919, 79
- Liu et al. (2022) Liu, J., Zhang, Q., & Qiu, K. 2022, Frontiers in Astronomy and Space Sciences, 9, 943556
- Lopez-Rodriguez et al. (2021) Lopez-Rodriguez, E., Beck, R., Clark, S. E., et al. 2021, ApJ, 923, 150
- Lopez-Rodriguez et al. (2022) Lopez-Rodriguez, E., Mao, S. A., Beck, R., et al. 2022, ApJ, 936, 92
- Loren et al. (1990) Loren, R. B., Wootten, A., & Wilking, B. A. 1990, ApJ, 365, 269
- Mackey & Lim (2011a) Mackey, J. & Lim, A. J. 2011a, MNRAS, 412, 2079
- Mackey & Lim (2011b) Mackey, J. & Lim, A. J. 2011b, MNRAS, 412, 2079
- Matthews et al. (2009) Matthews, B. C., McPhee, C. A., Fissel, L. M., & Curran, R. L. 2009, ApJS, 182, 143
- McKee & Ostriker (2007) McKee, C. F. & Ostriker, E. C. 2007, ARA&A, 45, 565
- Mestel (1966) Mestel, L. 1966, MNRAS, 133, 265
- Mestel & Spitzer (1956) Mestel, L. & Spitzer, L., J. 1956, MNRAS, 116, 503
- Mookerjea et al. (2018) Mookerjea, B., Sandell, G., Vacca, W., Chambers, E., & Güsten, R. 2018, A&A, 616, A31
- Motte et al. (1998) Motte, F., Andre, P., & Neri, R. 1998, A&A, 336, 150
- Mouschovias (2001) Mouschovias, T. 2001, in Astronomical Society of the Pacific Conference Series, Vol. 248, Magnetic Fields Across the Hertzsprung-Russell Diagram, ed. G. Mathys, S. K. Solanki, & D. T. Wickramasinghe, 515
- Mouschovias & Ciolek (1999) Mouschovias, T. C. & Ciolek, G. E. 1999, in NATO Advanced Study Institute (ASI) Series C, Vol. 540, The Origin of Stars and Planetary Systems, ed. C. J. Lada & N. D. Kylafis, 305
- Mouschovias et al. (2006) Mouschovias, T. C., Tassis, K., & Kunz, M. W. 2006, ApJ, 646, 1043
- Nakano & Nakamura (1978) Nakano, T. & Nakamura, T. 1978, PASJ, 30, 671
- Ngoc et al. (2023) Ngoc, N. B., Diep, P. N., Hoang, T., et al. 2023, ApJ, 953, 66
- Ngoc et al. (2021) Ngoc, N. B., Diep, P. N., Parsons, H., et al. 2021, ApJ, 908, 10
- Ortiz-León et al. (2017) Ortiz-León, G. N., Loinard, L., Kounkel, M. A., et al. 2017, ApJ, 834, 141
- Ostriker et al. (2001) Ostriker, E. C., Stone, J. M., & Gammie, C. F. 2001, ApJ, 546, 980
- Pabst et al. (2019) Pabst, C., Higgins, R., Goicoechea, J. R., et al. 2019, Nature, 565, 618
- Pabst et al. (2020) Pabst, C. H. M., Goicoechea, J. R., Teyssier, D., et al. 2020, A&A, 639, A2
- Pattle & Fissel (2019) Pattle, K. & Fissel, L. 2019, Frontiers in Astronomy and Space Sciences, 6, 15
- Pattle et al. (2023) Pattle, K., Fissel, L., Tahani, M., Liu, T., & Ntormousi, E. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 193
- Pattle et al. (2021) Pattle, K., Lai, S.-P., Di Francesco, J., et al. 2021, ApJ, 907, 88
- Pattle et al. (2022) Pattle, K., Lai, S.-P., Sadavoy, S., et al. 2022, MNRAS, 515, 1026
- Pattle et al. (2017) Pattle, K., Ward-Thompson, D., Berry, D., et al. 2017, ApJ, 846, 122
- Pattle et al. (2018) Pattle, K., Ward-Thompson, D., Hasegawa, T., et al. 2018, ApJ, 860, L6
- Pattle et al. (2015) Pattle, K., Ward-Thompson, D., Kirk, J. M., et al. 2015, MNRAS, 450, 1094
- Pillai et al. (2011) Pillai, T., Kauffmann, J., Wyrowski, F., et al. 2011, A&A, 530, A118
- Pillai et al. (2020) Pillai, T. G. S., Clemens, D. P., Reissl, S., et al. 2020, Nature Astronomy, 4, 1195
- Planck Collaboration et al. (2016) Planck Collaboration, Ade, P. A. R., Aghanim, N., et al. 2016, A&A, 586, A138
- Plummer (1911) Plummer, H. C. 1911, MNRAS, 71, 460
- Poglitsch et al. (2010) Poglitsch, A., Waelkens, C., Geis, N., et al. 2010, A&A, 518, L2
- Poidevin et al. (2013) Poidevin, F., Falceta-Gonçalves, D., Kowal, G., de Gouveia Dal Pino, E., & Mário Magalhães, A. 2013, ApJ, 777, 112
- Purcell (1979) Purcell, E. M. 1979, ApJ, 231, 404
- Santos et al. (2019) Santos, F. P., Chuss, D. T., Dowell, C. D., et al. 2019, ApJ, 882, 113
- Santos-Lima et al. (2012) Santos-Lima, R., Dal Pino, E. d. G., & Lazarian, A. 2012, The Astrophysical Journal, 747, 21
- Shu et al. (1987) Shu, F. H., Adams, F. C., & Lizano, S. 1987, ARA&A, 25, 23
- Skalidis et al. (2021) Skalidis, R., Sternberg, J., Beattie, J. R., Pavlidou, V., & Tassis, K. 2021, A&A, 656, A118
- Skalidis & Tassis (2021) Skalidis, R. & Tassis, K. 2021, A&A, 647, A186
- Soam et al. (2018) Soam, A., Maheswar, G., Lee, C. W., Neha, S., & Kim, K.-T. 2018, MNRAS, 476, 4782
- Soler (2019) Soler, J. D. 2019, A&A, 629, A96
- Soler et al. (2017) Soler, J. D., Ade, P. A. R., Angilè, F. E., et al. 2017, A&A, 603, A64
- Tram et al. (2023) Tram, L. N., Bonne, L., Hu, Y., et al. 2023, ApJ, 946, 8
- Tram et al. (2021) Tram, L. N., Hoang, T., Lee, H., et al. 2021, ApJ, 906, 115
- Tritsis et al. (2015) Tritsis, A., Panopoulou, G. V., Mouschovias, T. C., Tassis, K., & Pavlidou, V. 2015, MNRAS, 451, 4384
- Vaillancourt (2006) Vaillancourt, J. E. 2006, PASP, 118, 1340
- Ward-Thompson et al. (2023) Ward-Thompson, D., Karoly, J., Pattle, K., et al. 2023, ApJ, 946, 62
- Ward-Thompson et al. (2017) Ward-Thompson, D., Pattle, K., Bastien, P., et al. 2017, ApJ, 842, 66
- White et al. (2015) White, G. J., Drabek-Maunder, E., Rosolowsky, E., et al. 2015, MNRAS, 447, 1996
- Wilking et al. (2008) Wilking, B. A., Gagné, M., & Allen, L. E. 2008, in Handbook of Star Forming Regions, Volume II, ed. B. Reipurth, Vol. 5, 351
- Zinnecker & Yorke (2007) Zinnecker, H. & Yorke, H. W. 2007, ARA&A, 45, 481
- Zucker & Chen (2018) Zucker, C. & Chen, H. H.-H. 2018, ApJ, 864, 152
Appendix A Histogram of SOFIA/HAWC dataset
Figures 13 and 14 show the histogram of Stokes-, , SNRI, , , and SNRp of the SOFIA/HAWC data in band C and band D. Table 3 shows the summary of the mean and RMS of the distribution of Stokes-, , their uncertainties and SNRs.
Wavelength | (Jy/arcsec2) | (Jy/arcsec2) | SNRI | () | () | SNRP | ||||||
Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | |
89 m | 0.163 | 0.166 | 0.005 | 0.004 | 44.8 | 52.9 | 8.3 | 9.2 | 6.3 | 7.8 | 3.2 | 3.5 |
154 m | 0.143 | 0.142 | 0.001 | 0.001 | 177.5 | 214.1 | 5.7 | 4.1 | 1.2 | 1.5 | 10.3 | 9.9 |
Appendix B Magnetic field orientations and strengths at different wavelengths
In Figure 15, we compare the B-fields vectors obtained from HAWC data at 89 and 154 m (this study), and from POL-2 data at 850 m (Kwon et al. 2018). Vectors show the mean B-field position angles at the subregions in Oph-A.
Figure 16 shows the maps of B-field strength estimated using the ST method () at 89 and 154 m. Figure 17 shows the histogram distribution of the obtained with HAWC+ data at 89 and 154 m.
Appendix C Velocity dispersion traced by other molecular lines
To test whether the FWHM map of HCO+(43) line is a good choice, we further use the N2H+(32) molecular data observed by APEX to estimate the total velocity dispersion, . The majority of the Oph-A cloud exhibits the N2H+(32) emission showing 3 single Gaussian profiles; the middle peak is much stronger compared with the other two components. This strongest hyperfine line indicates the central velocity which is in the range of 2.54–3.97 km s-1, with an average velocity of 3.42 km s-1, in agreement with the systemic velocity of Oph-A of 3.44 km s-1 estimated by André et al. (2007). To each pixel we fit the spectrum with the sum of 3 single Gaussian profiles and then use the line width of the primary peak as the toward the computing pixel. We construct the map of the FWHM velocity dispersion, following the procedure presented in Section 3.2.2 for the case of the HCO+ molecular line.
Figure 18 shows the map of using the N2H+(32) molecular line toward Oph-A. The values ranges from 0.18 to 1.73 km s-1 with a median value of 0.95 km s-1, in good agreement with those obtained by HCO+(43) line. We note, however, that the spatial extent of the HCO+(43) emission line is large than that of the N2H+(32). Thus, we use the map of FWHM of the HCO+(43) line to calculate the strengths in Oph-A. Similarly, White et al. (2015) reported a mean FWHM velocity dispersion of 1.5 km s-1 using the C18O(32) data from JCMT/HARP in overall entire Ophiuchus cloud. Using the N2H+(10) data from IRAM, André et al. (2007) estimated an average value of the 1D nonthermal velocity dispersion toward Oph-A, =0.2 km s-1, corresponding to 0.47 km s-1. Observations of the NH3 hyperfine emission from the Green Bank Telescope (GBT) also returned a similar 1D velocity dispersion with the average value of 0.3 km s-1, corresponding to an average of FWHM of 0.65 km s-1 (Chen et al. 2019).
Appendix D Mass-to-flux ratio, Mach Alfvén Mach number, and beta plasma
Figure 19 show the maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C. Figure 20 show maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C (left panels) and D (right panels).
Appendix E relation using estimated with the ST method
Figure 21 shows the relation between and where B-field strength is estimated using the St method.
Appendix F Main structure of the ridge identification
We identify the elongated shape of the ridge in the densest region of Oph-A using the Python package RadFil. We first apply a threshold of 0.5 Jy/arcsec2 in the map of Stokes- at 154 m to build a mask for the main shape of the densest region (the ridge). The spine of the ridge is identified by using the Python package FilFinder777https://github.com/e-koch/FilFinder from Koch & Rosolowsky (2015). We then smooth the ridge spine and define a tangent line at every position on the ridge where the orientation of the ridge () at every position is identified. Figure 22 shows the schematic view of the ridge and the perpendicular cuts at each position on the ridge. The position angle of the tangent line at each position located on the ridge is defined as the orientation of the ridge at that position, measured east of north, and ranges from 0° to 180°.

### v2

@@
 Together with gravity, turbulence, and stellar feedback, magnetic fields (B-fields) are thought to play a critical role in the evolution of molecular clouds and star formation processes.
-Polarization of thermal dust emission is a popular tracer of B-fields in star-forming regions.
+The polarization of thermal dust emission is a popular tracer of B-fields in star-forming regions.
 Aims.
-We aim to map the morphology and measure the strength of B-fields of the nearby molecular cloud, rho Ophiuchus-A ( Oph-A), and then to understand the role of B-fields in regulating star formation and shaping the cloud.
+We aim to map the morphology and measure the strength of B-fields of the nearby molecular cloud, rho Ophiuchus-A ( Oph-A), to understand the role of B-fields in regulating star formation and in shaping the cloud.
 Methods.
-We have analyzed the far-infrared (FIR) polarization of thermal dust emission observed by SOFIA/HAWC at 89 and 154 m toward the densest part of Oph-A, which is irradiated by the nearby B3/4 star, Oph-S1.
+We analyzed the far-infrared (FIR) polarization of thermal dust emission observed by SOFIA/HAWC at 89 and 154 m toward the densest part of Oph-A, which is irradiated by the nearby B3/4 star, Oph-S1.
 These FIR polarimetric maps cover an area of (corresponding to pc2) with an angular resolution of 7.8 and 13.6, respectively.
 Results.
-The Oph-A cloud exhibits well-ordered B-fields with magnetic orientations mainly perpendicular to the ridge of the cloud toward the densest region.
-We obtain a map of B-field strengths in the range of 0.2-2.5 mG, using the Davis–Chandrasekhar-Fermi method.
-The B-fields are strongest at the densest part of the cloud, which is associated with the starless core SM1, and decreases toward the outskirts of the cloud.
+The Oph-A cloud exhibits well-ordered B-fields with magnetic orientations that are mainly perpendicular to the ridge of the cloud toward the densest region.
+We obtained a map of B-field strengths in the range of 0.2-2.5 mG, using the Davis-Chandrasekhar-Fermi (DCF) method.
+The B-fields are strongest at the densest part of the cloud, which is associated with the starless core SM1, and then decrease toward the outskirts of the cloud.
 By calculating the map of the mass-to-flux ratio, Alfvén Mach number, and plasma parameter in Oph-A, we find that the cloud is predominantly magnetically sub-critical, sub-Alfvénic, which implies that the cloud is supported by strong B-fields that dominate over gravity, turbulence, and thermal gas energy.
-Measured B-field strengths at two densest subsregions using other methods that account for the compressible mode are relatively lower than that measured with the DCF method but do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation.
-A virial analysis suggests that the cloud is gravitationally unbound, which is consistent with the previous detection of numerous starless cores in the cloud.
-Comparing the magnetic pressure with the radiation pressure from the Oph-S1 star, we find that B-fields are sufficiently strong to support the cloud against radiative feedback and to regulate the shape of the cloud.
+The measured B-field strengths at the two densest subsregions using other methods that account for the compressible mode are relatively lower than that measured with the DCF method.
+However, these results do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation.
+Our virial analysis suggests that the cloud is gravitationally unbound, which is consistent with the previous detection of numerous starless cores in the cloud.
+By comparing the magnetic pressure with the radiation pressure from the Oph-S1 star, we find that B-fields are sufficiently strong to support the cloud against radiative feedback and to regulate the shape of the cloud.
 Key Words.:
@@
 B-fields are thought to play a critical role in the star formation process by supporting the material in the cloud against self-gravitational collapse (see, e.g., Mouschovias, 2001; Crutcher, 2012; Pattle & Fissel, 2019; Pattle et al., 2023).
-The collapse can occur when the dense core is magnetically super-critical, with the ratio of mass to magnetic flux (known as mass-to-flux ratio) greater than one, i.e., when B-fields are not strong enough to prevent the gravitational collapse.
+The collapse can occur when the dense core is magnetically super-critical, with the ratio of mass to magnetic flux (known as mass-to-flux ratio) greater than 1; that is, when B-fields are not strong enough to prevent the gravitational collapse.
 The collapse of super-critical cores will lead to the formation of protostars.
 Conversely, when the core is sub-critical, strong B-fields can act against gravitational collapse.
-However, several mechanisms might enhance the mass-to-flux ratio and enable the core to collapse, including ambipolar diffusion (see e.g., Mestel & Spitzer, 1956; Mouschovias et al., 2006; Liu et al., 2022) and magnetic reconnection diffusion (Lazarian & Vishniac, 1999; Santos-Lima et al., 2012).
+However, several mechanisms might enhance the mass-to-flux ratio and cause the core to collapse, including ambipolar diffusion (see e.g., Mestel & Spitzer, 1956; Mouschovias et al., 2006; Liu et al., 2022) and magnetic reconnection diffusion (Lazarian & Vishniac, 1999; Santos-Lima et al., 2012).
 Feedback from massive stars (O- or B-type) can influence the properties of the parental molecular cloud and the star formation processes (e.g., Krumholz et al., 2014; Pabst et al., 2019, 2020).
-Three-dimensional (3D) magnetohydrodynamic (MHD) simulations have suggested that B-fields also influence the structure of clouds along with radiation pressure feedback from massive stars (e.g., Henney et al., 2009; Mackey & Lim, 2011a).
+Three-dimensional magnetohydrodynamic (3D MHD) simulations have suggested that B-fields also influence the structure of clouds along with radiation pressure feedback from massive stars (e.g., Henney et al., 2009; Mackey & Lim, 2011a).
 It has been shown that when the B-field orientations are parallel to the ionizing radiation field, the cloud will tend to have a flattened shape, while the cloud will have a column-like or pillar structure if the B-fields are perpendicular to the radiation field (Mackey & Lim, 2011b).
 These findings from simulations have been in agreement with observations and could explain various globular and elongated substructures in the IC 1396 region (Soam et al., 2018), the famous pillar structure in the M16 region (Pattle et al., 2018), and the Horsehead Nebula (Hwang et al., 2023).
-Interstellar dust grains are known to have aspherical shapes and can be systematically aligned with a fixed direction in space, e.g., interstellar B-fields (Hall, 1949; Hiltner, 1949).
+Interstellar dust grains are known to have aspherical shapes and can be systematically aligned with a fixed direction in space, for instance, interstellar B-fields (Hall, 1949; Hiltner, 1949).
 The alignment of interstellar dust grains with B-fields involves the internal alignment of the grain axis of the maximum inertia moment (also the shortest axis) with its angular momentum (i.e., internal alignment) and the alignment of the angular momentum with B-fields (external alignment).
 Internal alignment is caused by internal relaxation processes, including Barnett relaxation and inelastic relaxation (Purcell, 1979).
-The external alignment is caused by the Radiative Torque (RAT) alignment mechanism (Draine & Weingartner, 1997; Lazarian & Hoang, 2007) and by the more comprehensive magnetically-enhanced RAT mechanism (MRAT, Hoang & Lazarian, 2016).
+The external alignment is caused by the radiative torque (RAT) alignment mechanism (Draine & Weingartner, 1997; Lazarian & Hoang, 2007) and by the more comprehensive magnetically enhanced RAT mechanism (MRAT, Hoang & Lazarian, 2016).
 The internal and external alignment processes establish the alignment of the grain with its minor axis parallel to the B-fields (see, e.g., Hoang et al., 2022).
 Dust grains absorb or scatter starlight in the ultraviolet (UV), optical, or near-infrared (NIR) range.
-However, aligned dust grains tend to absorb/scatter the electric field component of the starlight along their major axis more efficiently.
+However, aligned dust grains tend to absorb and scatter the electric field component of the starlight along their major axis more efficiently.
 This results in the polarization of background starlight with the polarization vector parallel to the grain’s minor axis.
 Thus, the polarization of the starlight seen in the UV to optical or NIR wavelength has the polarization orientation parallel to the B-field orientation.
-Aligned dust grains also re-emit the energy absorbed from starlight in the longer wavelength from far-infrared (FIR) to sub-millimeter (sub-mm), so-called thermal dust emission.
+Aligned dust grains also re-emit the energy absorbed from starlight in the longer wavelength from far-infrared (FIR) to sub-millimeter (sub-mm), the so-called thermal dust emission.
 As the electric field component of starlight is absorbed more along the grain’s major axis, the dust grain will re-emit more energy with the electric field component along this major axis than along the minor axis.
@@
 Thus, if the position angle of the polarized thermal dust emission is known, one can rotate the polarization angle by 90° to infer the position angle of the B-fields in the medium.
-Linear polarized thermal dust emission has been routinely used to trace B-fields in various environments of molecular clouds in the Milky Way (see, e.g., Ward-Thompson et al., 2017; Chuss et al., 2019; Pattle et al., 2022) and extra-galactic regions (see, e.g., Lopez-Rodriguez et al., 2021, 2022; Borlaff et al., 2023).
-The Davis-Chandrasekhar-Fermi method (DCF; Davis, 1951; Chandrasekhar & Fermi, 1953) has been widely used to measure the strengths of the B-field projected in the plane-of-the-sky, , based on the assumption that the kinetic energy of turbulent gas balances to that of turbulent B-fields.
-Adopting the mean value of B-field strengths along the line-of-sight () measured using the Zeeman technique over different molecular clouds, Crutcher et al.
+Linear polarized thermal dust emission has been routinely used to trace B-fields in various environments of molecular clouds in the Milky Way (see, e.g., Ward-Thompson et al., 2017; Chuss et al., 2019; Pattle et al., 2022) and extragalactic regions (see, e.g., Lopez-Rodriguez et al., 2021, 2022; Borlaff et al., 2023).
+The Davis-Chandrasekhar-Fermi method (DCF; Davis, 1951; Chandrasekhar & Fermi, 1953) has been widely used to measure the strengths of the B-field projected on the plane of the sky, , based on the assumption that the kinetic energy of turbulent gas balances to that of turbulent B-fields.
+Adopting the mean value of B-field strengths along the line of sight () measured using the Zeeman technique over different molecular clouds, Crutcher et al.
 (2010) derived a relationship between the B-field strengths and gas volume densities ().
@@
 Pattle et al.
-(2023) compiled a large number of measurements of the B-field strength projected in the plane-of-the-sky () estimated using the DCF method and showed that these generally follow Crutcher et al.
-(2010)’s results until cm-3.
-Beyond this value, the distribution is more sparse (see Figure 2a in Pattle et al., 2023).
+(2023) compiled a large number of measurements of the B-field strength projected on the plane of the sky () estimated using the DCF method.
+These authors showed that these generally follow the results from Crutcher et al.
+(2010) until cm-3.
+Beyond this value, the distribution is more sparse (see Fig.
+2a in Pattle et al., 2023).
 Theoretical studies have suggested that the power-law index varies between 2/3 for weak B-fields (Mestel, 1966) and 0.5 for strong B-fields (Mouschovias & Ciolek, 1999).
 Therefore, the maps of the morphology and strength of the B-fields within a specific cloud are essential to understanding the role of the local B-fields on star formation within the cloud.
-Previous authors attempted to derive the map of B-fields toward some well-known star-forming regions, for instance, OMC-1 (Guerra et al.
-2021, Hwang et al.
-2021), 30 Doradus (30 Dor, Tram et al.
-2023), and the G11.11–0.12 filament (Ngoc et al.
-2023).
-These studies revealed significant variations in the B-field morphology and strength across the clouds and demonstrated that B-fields strongly influence the cloud’s evolution and, thus, the star formation activities.
+Previous authors have attempted to derive the map of B-fields toward some well-known star-forming regions, for instance, OMC-1 (Guerra et al., 2021; Hwang et al., 2021), 30 Doradus (30 Dor, Tram et al., 2023), and the G11.11-0.12 filament (Ngoc et al., 2023).
+These studies revealed significant variations in the B-field morphology and strength across the clouds, demonstrating that B-fields strongly influence the cloud’s evolution and, thus, star formation activities.
 In this work, we focus on a specific system of the Ophiuchus-A (hereafter Oph-A) molecular cloud to study the effects of stellar feedback and B-fields on the evolution of the cloud and star formation.
@@
 The Oph-A cloud was initially identified as a dense gas clump based on DCO+ observations performed by Loren et al.
-(1990) among other clumps named from Oph-B to Oph-F.
-This region has high extinctions (20 mag) corresponding to the high gas column densities compared to other regions, which are typically more quiescent (see Figure 1).
+(1990), along with other clumps named Oph-B to Oph-F.
+This region has high extinctions (20 mag) corresponding to the high gas column densities compared to other regions, which are typically more quiescent (see Fig.
+1).
 Being one of the closest low-mass star-forming regions in our Galaxy, located at a distance of 137 pc (Ortiz-León et al., 2017), the gas and dust content in the Oph–A region has been studied extensively at almost every wavelength by both ground-based and space instruments for many years (see e.g., Wilking et al., 2008, for a review).
@@
 The Oph-A cloud is mostly influenced by a nearby embedded early high-mass B3/4 star (Brown & Zuckerman, 1975; Mookerjea et al., 2018), Oph-S1, which has a mass of 8 (Hamaguchi et al., 2003).
-The Oph-S1 star is surrounded by a compact H ii region (Andre et al., 1988) and is believed to trigger the star formation activity in Oph–A, especially along the ridge of the dense cloud with indications of protostellar clumps (see Figure 2a in Motte et al., 1998).
-The star formation activity in the Oph-A is also believed to be affected by the Sco OB2 association located at the western part of the cloud at a distance of 145 pc (de Zeeuw et al., 1999).
-B-fields in Oph-A were investigated using data observed in various wavelengths: the NIR (JHK bands) using SIRPOL on the Infrared Survey Facility (IRSF) 1.4–m telescope (Kwon et al., 2015), FIR by the on the 2.7–m Stratospheric Observatory For Infrared Astronomy (SOFIA)/HAWC (Santos et al., 2019), sub-mm by the 15–m James Clerk Maxwell Telescope (JCMT)/POL–2 (Kwon et al., 2018).
+The Oph-S1 star is surrounded by a compact H ii region (Andre et al., 1988) and is believed to trigger the star formation activity in Oph–A, especially along the ridge of the dense cloud with indications of protostellar clumps (see Fig.
+2a in Motte et al., 1998).
+The star formation activity in the Oph-A is also believed to be affected by the Sco OB2 association located in the western part of the cloud at a distance of 145 pc (de Zeeuw et al., 1999).
+We investigated B-fields in Oph-A using data observed in various wavelengths: NIR (JHK bands) using imaging polarimeter SIRPOL (Kandori et al., 2006) mounted on the 1.4-m Infrared Survey Facility (IRSF) telescope (Kwon et al., 2015), FIR by the High-resolution Airborne Wideband Camera-plus (HAWC; Harper et al., 2018) on board the 2.7-m Stratospheric Observatory For Infrared Astronomy (SOFIA) (Santos et al., 2019), and sub-mm using the linear polarimeter POL-2 placed in front of the Submillimetre Common-User Bolometer Array 2 (SCUBA-2; Holland et al., 2013) camera mounted on the 15-m James Clerk Maxwell Telescope (JCMT) (Kwon et al., 2018).
 Santos et al.
@@
 Kwon et al.
-(2018) estimated the values in Oph-A using the JCMT/POL–2 data at 850 m toward a few densest parts of the cloud, yielding values in the range of 0.5–5 mG.
-The highest strength was measured toward the densest part of the cloud associated with the starless core SM1 (see Figure 1).
+(2018) estimated the values in Oph-A, using the JCMT/POL-2 data at 850 m, toward a few of the densest parts of the cloud, yielding values in the range of 0.5–5 mG.
+The highest strength was measured toward the densest part of the cloud associated with the starless core SM1 (see Fig.
+1).
 These strengths in Oph-A are comparable to values found in the OMC region (Hwang et al., 2021).
@@
 The Oph-A region provides an ideal isolated system for studying the impact of the B-fields in the star-forming region and investigating the interplay between B-fields and the stellar feedback into the molecular cloud.
-In this paper, we aim to:
-1.
-determine the B-field morphology and strengths in the entire Oph-A region using the SOFIA/HAWC dust polarization data;
+In this paper, we aim to: 1) determine the B-field morphology and strengths in the entire Oph-A region using the SOFIA/HAWC dust polarization data; 2) investigate the role of B-fields relative to gravity and turbulence in the star formation activity in the cloud; and 3) quantify the combined effect of B-fields and radiation pressure from the nearby Oph-S1 star on the shape of the Oph-A cloud.
+The paper is organized as follows.
+We describe a set of archival observational data in Sect.
 2.
-investigate the role of B-fields relative to gravity and turbulence in the star formation activity in the cloud;
-3.
-quantify the combined effect of B-fields and radiation pressure from the nearby Oph-S1 star on the shape of the Oph-A cloud.
-The paper is organized as follows.
-We describe a set of archival observational data in Section 2.
-We present the results of mapping the B-field morphology and strength using various techniques in Section 3 and discuss the roles of B-fields relative to gravity, turbulence, and radiation pressure feedback from the Oph-S1 star in Section 4.
-Finally, we summarize our results in Section 5.
+We present the results of mapping the B-field morphology and strength using various techniques in Sect.
+3 and discuss the roles of B-fields relative to gravity, turbulence, and radiation pressure feedback from the Oph-S1 star in Sect.
+4.
+Finally, we summarize our results in Sect.
+5.
 2 Observational data
 2.1 Dust polarization data from SOFIA/HAWC
-We use the FIR polarimetric observations toward Oph-A from the High-resolution Airborne Wideband Camera-plus (HAWC; Harper et al., 2018) onboard SOFIA.
+We use the FIR polarimetric observations from the onboard SOFIA/HAWC toward Oph-A.
 The observations were done on 2017 May 17, under the proposal ID (PI: Darren Dowell), using the nod-match-chop mode, with 154° chop angle, 480″ chop throw, and 4-points dithers.
-Here, we use the SOFIA/HAWC Level–4 data for two bands C and D, centered at 89 and 154 m, respectively, which is fully calibrated and can be directly downloaded from the SOFIA archive111https://irsa.ipac.caltech.edu/Missions/sofia.html.
-The final polarimetric maps cover an area of toward the central region in Oph-A with the angular resolutions of 7.8 and 13.6 full-width-half-maximum (FWHM) at 89 and 154 m, respectively.
+Here, we use the SOFIA/HAWC Level-4 data for two bands C and D, centered at 89 and 154 m, respectively, which is fully calibrated and can be directly downloaded from the SOFIA archive111https://irsa.ipac.caltech.edu/Missions/sofia.html.
+The final polarimetric maps cover an area of (corresponding to pc2) toward the central region in Oph-A with the angular resolutions of 7.8 and 13.6 full-width-half-maximum (FWHM) at 89 and 154 m, respectively.
 The pixel size of HAWC maps is 2.0″ at 89 m and 3.4″ at 154 m.
@@
 (2019).
-From the maps of the Stokes-, Stokes-, Stokes- and their uncertainties , , , the maps of polarization fraction () and its corresponding uncertainty () are calculated.
+From the maps of the Stokes-, Stokes-, and Stokes- (along with their uncertainties , , ), the maps of polarization fraction () and its corresponding uncertainty () were calculated.
 The de-biased polarization fraction, , is calculated as: (Vaillancourt, 2006).
 The maps of polarization angle () and its corresponding uncertainty () in degrees can be derived from Stokes- and Stokes-.
-We refer the reader to Gordon et al.
+We refer to Gordon et al.
 (2018) for a full description of estimates of the polarization fraction, polarization angle, and their corresponding uncertainties.
@@
 We mask out from SOFIA/HAWC maps the pixels for which ¡54 for band C and ¡214 for band D.
-In both bands, we further remove the pixels for which ¡3 and ¿50% from HAWC maps (see Appendix A).
+In both bands, we further removed the pixels for which we have ¡3 and ¿50% from the HAWC maps (see Appendix A).
 2.2 Spectroscopic data
-We use the spectroscopic observations of HCO+(43) line at 356.734 GHz obtained from the Heterodyne Array Receiver Program (HARP; Buckle et al., 2009) mounted on JCMT.
-The observations were done during June and July 2011 (Proposal ID: M11AU13), covering the Oph-A region with a beam size (or FWHM) of 14.
-We use the reduced spectral map of HCO+(43) line directly downloaded from the JCMT science archive222https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/.
-The data cube has an original spectral resolution of 0.025 km s-1 and is smoothed to have a final spectral resolution of 0.5 km s-1.
-For comparison to the HCO+(43) data, we also use observations of N2H+(32) line at 279.515 GHz obtained from the APEX-2 instrument on the 12-m Atacama Pathfinder EXperiment telescope (APEX; Güsten et al., 2006).
+We used the spectroscopic observations of HCO+(43) line at 356.734 GHz obtained from the Heterodyne Array Receiver Program (HARP; Buckle et al., 2009) mounted on JCMT.
+The observations were carried out during June and July 2011 (Proposal ID: M11AU13), covering the Oph-A region with a beam size (or FWHM) of 14.
+We used the reduced spectral map of HCO+(43) line directly downloaded from the JCMT science archive222https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/.
+The data cube had an original spectral resolution of 0.025 km s-1 and was smoothed to have a final spectral resolution of 0.5 km s-1.
+For comparison to the HCO+(43) data, we also used observations of N2H+(32) line at 279.515 GHz obtained from the APEX-2 instrument on the 12-m Atacama Pathfinder EXperiment telescope (APEX; Güsten et al., 2006).
 The data cube has an original spectral resolution of 0.082 km s-1.
-Here, we use the back-end data cube with a spectral resolution of 0.25 km s-1, covering the Oph-A region with the beam size of 22 (for details, see Liseau et al., 2015).
+Here, we used the back-end data cube with a spectral resolution of 0.25 km s-1, covering the Oph-A region with the beam size of 22 (for details, see Liseau et al., 2015).
 2.3 Dust temperature and column density maps
-We adopt the maps of dust temperature () and gas column density () from Santos et al.
+We adopted the maps of dust temperature () and gas column density () from Santos et al.
 (2019).
-These maps were obtained from the fitting of a modified black body to the spectral energy distribution (SED) using the continuum flux maps from Herschel/PACS (Poglitsch et al., 2010) at 70, 100, and 160 m and have the FWHM of 11.4 (for details, see Section 3.2 in Santos et al., 2019).
-Figure 1 shows the spatial distribution toward Oph-A, overlaid with contours of column density .
-In this study, we will use the map of as the map of the gas kinetic temperature, assuming that the gas is in the local thermodynamic equilibrium (LTE) and well-coupled with the dust.
+These maps were obtained from the fitting of a modified black body to the spectral energy distribution (SED) using the continuum flux maps from Herschel/PACS (Poglitsch et al., 2010) at 70, 100, and 160 m, with an FWHM of 11.4 (for details, see Sect.
+3.2 in Santos et al., 2019).
+Figure 1 shows the spatial distribution toward Oph-A, overlaid with contours of column density, .
+In this study, we use the map of as the map of the gas kinetic temperature, assuming that the gas is in the local thermodynamic equilibrium (LTE) and well-coupled with the dust.
 3 Magnetic fields in Oph-A
@@
 Here, the position angle of B-fields () is measured east of north, and in the range between 0 and 180.
-The polarimetric data at both wavelengths can well reveal B-fields toward the dense region in Oph-A where Stokes-¿0.12 Jy/arcsec2 (the area inside the lowest contour in each map).
+The polarimetric data at both wavelengths can well reveal B-fields toward the dense region in Oph-A, where Stokes-¿0.12 Jy/arcsec2 (the area inside the lowest contour in each map).
 The data at 89 m can further resolve the B-fields morphology extended toward the warmer dust region located at the east side of the Oph-S1 star.
-The B-field vectors in Oph-A are generally well-ordered and largely follow the material structure in low- to intermediate-densities regimes (see Figure 2).
-B-field orientations seem to be bent toward the higher-density region in the center of the map and are mostly perpendicular to the ridge of the cloud.
-On the eastern side of the map, the orientation of B-field vectors tends to vary in the counter-clockwise direction from North to South, i.e., the position angle of B-fields increases.
-Particularly, in the North-East part of the map, B-field vectors run along the North-East to South-West direction with a median value of 56° and 49° at 89 and 154 m, respectively.
-This feature is also seen in a small area in the North-West part of the map.
+The B-field vectors in Oph-A are generally well-ordered and largely follow the material structure in low- to intermediate-densities regimes (see Fig.
+2).
+B-field orientations seem to be bent toward the higher density region in the center of the map and are mostly perpendicular to the ridge of the cloud.
+On the eastern side of the map, the orientation of B-field vectors tends to vary in the counter-clockwise direction from north to south; thus, the position angle of B-fields increases.
+Particularly, in the northeast part of the map, B-field vectors run along the northeast to southwest direction with a median value of 56° and 49° at 89 and 154 m, respectively.
+This feature is also seen in a small area in the northwest part of the map.
 Toward the east side, B-field vectors orient nearly to the east-west direction, with median values of 83° at 89 m and 75° at 154 m.
-On the other hand, B-field vectors in the South-East region tend to orient consistently to the South-East to North-West directions, with median values of 92° and 85° at 89 and 154 m, respectively.
-Interestingly, the data at 154 m shows B-field vectors running nearly vertically, i.e., along the North-South direction toward a small area at the North part of the map, which is not seen in the data at 89 m (see Figure 2 and also in Figure 15).
-In addition, there are two outstanding components of B-fields toward the regions characterized by the highest H2 column densities: one component associated with the peak of continuum emission at 154 m runs along North-East to South-West direction with the median B-field orientation of 70; the other is located at the North of the first component and associated with the peak of continuum emission at 89 m, which exhibits the B-field vectors along the East-West direction with a median position angle of 88° and 81° at 89 and 154 m, respectively (see Figure 15).
-We will further discuss the B-field morphology together with B-field strengths in Section 4.1.
+On the other hand, B-field vectors in the southeast region tend to orient consistently to the southeast to northwest directions, with median values of 92° and 85° at 89 and 154 m, respectively.
+Interestingly, the data at 154 m shows B-field vectors running nearly vertically, that is, along the north-south direction toward a small area at the north part of the map, which is not seen in the data at 89 m (see Fig.
+2 and also Fig.
+15).
+In addition, there are two outstanding components of B-fields toward the regions characterized by the highest H2 column densities: one component associated with the peak of continuum emission at 154 m runs along the northeast to southwest direction with the median B-field orientation of 70; the other is located at the north of the first component and associated with the peak of continuum emission at 89 m, which exhibits the B-field vectors along the east-west direction with a median position angle of 88° and 81° at 89 and 154 m, respectively (see Fig.
+15).
+We will further discuss the B-field morphology together with B-field strengths in Sect.
+4.1.
 3.2 Measurement of magnetic field strength using the DCF method
-We first use the classical DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953) (hereafter we refer to as the DCF method) to calculate the strengths of the B-field projected in the plane-of-the-sky () in the entire region of the Oph-A cloud.
-We will compare our obtained results with those obtained using more advanced techniques in Section 3.3.
+We first used the classical DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953) (hereafter we refer to as the DCF method) to calculate the strengths of the B-field projected in the plane-of-the-sky () in the entire region of the Oph-A cloud.
+We compared our obtained results with those obtained using the more advanced techniques described in Sect.
+3.3.
 According to the DCF method, the B-field strength can be estimated by Ostriker et al. (2001):
-where is the correction factor, is the gas mass density in units of g cm-3, =2.8 is the mean molecular weight per hydrogen molecule (Kauffmann et al., 2008), is the mass of the hydrogen atom, is the gas volume density in units of cm-3), is the one-dimensional (1D) non-thermal velocity dispersion along the line-of-sight of the ideal gas species that traces the dense region as similar as by dust polarized emission in units of km s-1, and is the polarization angle dispersion in degrees.
-The correction factor is uncertain, and numerical simulations by Ostriker et al.
+where is the correction factor, is the gas mass density in units of g cm-3, =2.8 is the mean molecular weight per hydrogen molecule (Kauffmann et al., 2008), and is the mass of the hydrogen atom, is the gas volume density in units of cm-3.
+Then, is the one-dimensional (1D) non-thermal velocity dispersion along the line of sight (LOS) of the ideal gas species that traces the dense region, similarly to the dust-polarized emission in units of km s-1, and is the polarization angle dispersion in degrees.
+The correction factor, , is uncertain, and numerical simulations by Ostriker et al.
 (2001) show when the polarization angle dispersion does not exceed 25.
-However, varying values of will result in different strengths (see e.g., Skalidis et al., 2021; Liu et al., 2021, for detail of discussions).
-Here, we use the formula as appeared in Crutcher (2004):
+However, varying values of will result in different strengths (see e.g., Skalidis et al., 2021; Liu et al., 2021, for more detailed discussions).
+Here, we use the formula given in Crutcher (2004):
 where is non-thermal FWHM velocity dispersion measured in units of km s-1, and the correction factor has been used. This formula is widely used to estimate via the DCF technique.
@@
 It is essential to see how the B-field orientation changes locally within the molecular cloud.
-To do that, we calculate the dispersion in the polarization angle distribution () across the Oph-A region.
-We use the method presented in Hwang et al.
-(2021) for which is calculated pixel-by-pixel throughout the map of Oph-A.
+To do so, we calculated across the Oph-A region.
+We used the method presented in Hwang et al.
+(2021), for which is calculated pixel-by-pixel throughout the map of Oph-A.
 We summarize the procedure of calculating toward the pixel in the map as follows:
 1.
-We work with a box covering the pixel with a size of 99 pixels, corresponding to 2 beam sizes of the SOFIA/HAWC observations (i.e., 15.6 and 36.4 in bands C and D, respectively).
+We used a box covering the pixel with a size of 99 pixels, corresponding to two beam sizes of the SOFIA/HAWC observations (i.e., 15.6 and 36.4 in bands C and D, respectively).
 2.
-We calculate the mean value of polarization angles within the small box, , which implies the local mean polarization angle toward the pixel .
+We calculated the mean value of polarization angles within the small box, , which implies the local mean polarization angle toward the pixel.
 3.
-For each pixel within the small box, we calculate the difference between the observed polarization angle and the mean value.
-We compute the root mean square (RMS) of the distribution of angle differences in the box.
+For each pixel within the small box, we calculated the difference between the observed polarization angle and the mean value.
+We computed the root mean square (RMS) of the distribution of angle differences in the box.
 This RMS value indicates the polarization angle dispersion calculated toward the pixel.
@@
 The uncertainty of the polarization angle dispersion toward the pixel is propagated from the uncertainty of polarization angles in the small box and determined as the mean value of polarization angle uncertainties within the small box.
-By repeating the process for all pixels in the entire region, we obtain the map of in Oph-A.
+By repeating the process for all pixels in the entire region, we obtained the map of in Oph-A.
 We note that if there are more than 50 of pixels undefined (i.e., having NaN value of ) in the computing box, the result of polarization angle dispersion calculation toward pixel is ignored.
-We apply this criterion to keep the high quality of the estimate.
+We applied this criterion to keep the high quality of the estimate.
 It is also necessary to de-bias the polarization angle dispersion due to the uncertainty of the observed polarized angles.
 Figure 3 shows the maps of polarization angle dispersion toward Oph-A using HAWC data in two bands C (left panel) and D (right panel), respectively.
-The polarization angle dispersion calculated using data in bands C and D are in ranges 2.3–32.2 and 1.3–31.5, which correspond to the median values of 5.7° and 5.8°, respectively.
+The polarization angle dispersion calculated using data in bands C and D cover ranges of 2.3–32.2 and 1.3–31.5, which correspond to the median values of 5.7° and 5.8°, respectively.
 The differences of in two wavelengths imply that the data at the shorter wavelength tends to trace warm dust surrounding the high-mass star S1 and toward the outer layers; whereas the data at the longer wavelength probes the colder dust in the denser layers of the cloud.
 We note, however, that in maps of in both bands, there are a few pixels where (27 pixels in band C and 45 pixels in band D).
-Thus, we do not use the value in these pixels to estimate in Section 3.2.4.
+Thus, we do not use the value in these pixels to estimate in Sect.
+3.2.4.
 3.2.2 Velocity dispersion map
-We use the HCO+(43) molecular data observed by JCMT/HARP to estimate the total velocity dispersion, .
-This molecular line has a critical density of 106 cm-3 (Greve et al., 2009) and thus is considered a good tracer to probe the dense region in the Oph-A cloud.
+We used the HCO+(43) molecular data observed by JCMT/HARP to estimate the total velocity dispersion, .
+This molecular line has a critical density of 106 cm-3 (Greve et al., 2009) and thus is considered a good tracer for probing the dense region in the Oph-A cloud.
 The HCO+(43) emission is well-represented by a single Gaussian profile.
@@
 The central velocity () across Oph-A covers the range from 2.25 to 3.76 km s-1 with a median value of 3.12 km s-1, which is in agreement with the systemic velocity of Oph-A of 3.44 km s-1 (André et al., 2007).
-We use the line width of the Gaussian as the measure of total velocity dispersion toward each pixel.
-To construct the map of the 1D non-thermal (turbulent) velocity dispersion (), we subtract the thermal component () from the map of the total velocity dispersion, given by:
+We used the line width of the Gaussian as the measure of total velocity dispersion toward each pixel.
+To construct the map of the 1D non-thermal (turbulent) velocity dispersion (), we subtracted the thermal component () from the map of the total velocity dispersion, given by:
 where =29 amu is the mass of the HCO+ molecule, is the Boltzmann constant, and is the kinetic temperature of the gas.
@@
 Thus, the gas kinetic temperature is equal to the excitation temperature in all levels () and equal to the dust temperature: .
-We convolve the map of to the lower resolution of the molecular line map (14) to obtain the same pixel size on both maps.
-We convert the velocity dispersion of the non-thermal component into the FWHM velocity dispersion, given as .
+We convolved the map of to the lower resolution of the molecular line map (14) to obtain the same pixel size on both maps.
+We converted the velocity dispersion of the non-thermal component into the FWHM velocity dispersion, given as .
 The uncertainty of non-thermal FWHMs is propagated from the uncertainty of the line width obtained from the Gaussian fit.
@@
 The values are in the range from 0.62 to 2.98 km s-1 with a median value of 1.19 km s-1.
-We find a good agreement between our result obtained with HCO+(43) and those from other line tracers, i.e., 0.95 km s-1 using N2H+(32) from APEX and 1.5 km s-1 using C18O(32) from JCMT/HARP (see Appendix C for details).
+We find a good agreement between our result obtained with HCO+(43) and those from other line tracers, namely, 0.95 km s-1 using N2H+(32) from APEX and 1.5 km s-1 using C18O(32) from JCMT/HARP (see Appendix C for details).
 Thus, it is plausible to use results from the HCO+(43) line in the following analysis.
 3.2.3 Gas volume density map
-We estimate the gas volume, , across the dense cloud from .
+We estimated the gas volume, , across the dense cloud from .
 Here, we assume that the cloud has a spherical Plummer-like shape (Liseau et al., 2015; Santos et al., 2019), which is frequently used to describe the globular dense cores (Plummer, 1911).
 We consider a sphere centered at the peak position of , located at (RA, Dec)= (, ), which is also associated with the position of the starless core SM1 (the separation is much less than 1″).
-We take the radius of the cloud (0.074 pc, Santos et al., 2019).
+We took the radius of the cloud (0.074 pc, Santos et al., 2019).
 To describe the effect of stellar feedback from the high-mass Oph-S1 star, we further add a spherical shell with a radius of surrounding this star.
-We also assume that the positions of peak column density and the S1 star are located in the same plane perpendicular to the line of sight.
-The top panels in Figure 5 show the 3D schematic view of the cloud (left) and a cross-section on the -plane at a given position (right).
-The - and -axes perpendicular to each other are in the plane of the sky and the axis is along the line of sight.
+We also assume that the positions of peak column density and the S1 star are located in the same plane perpendicular to the LOS.
+The top panels in Fig.
+5 show the 3D schematic view of the cloud (left) and a cross-section on the -plane at a given position (right).
+The - and -axes perpendicular to each other are in the plane of the sky and the axis is along the LOS.
 The center of the sphere is located at .
 The depth of the cloud at each position () along axis is calculated by:
-where and are the radius of horizontal circles perpendicular to the plane of sky at position (see top right panel in Figure 5).
-We obtain the map of , shown in the bottom left panel in Figure 5.
+where and are the radius of horizontal circles perpendicular to the plane of sky at position (see top right panel in Fig.
+5).
+We obtain the map of , shown in the bottom left panel in Fig.
+5.
 The values of are in the range of .
 We then infer from using the formula:
-The bottom right panel in Figure 5 shows the spatial extent of across Oph-A, overlaid by the contours representing the continuum emission in band D (Stokes-).
+The bottom right panel in Fig.
+5 shows the spatial extent of across Oph-A, overlaid by the contours representing the continuum emission in band D (Stokes-).
 The map has an angular resolution of 11.4″and a pixel size of 3.2, similar to that of the map.
@@
 3.2.4 Magnetic field strength maps
-We use maps of polarization angle dispersion (), non-thermal FWHM velocity dispersion (), and gas volume density () to measure in the entire region of Oph-A.
-All input maps are smoothed to the angular resolution of 14, corresponding to that of the map.
-We then reconstruct the maps of at 89 and 154 m toward Oph-A using Eq.
-2, which are shown in Figure 6.
+We used the maps of , , and to measure in the entire region of Oph-A.
+All input maps were smoothed to the angular resolution of 14, corresponding to that of the map.
+We then reconstructed the maps of at 89 and 154 m toward Oph-A using Eq.
+2, which are shown in Fig.
+6.
 The values vary from mG and from mG for data at 89 and 154 m, respectively.
@@
 In addition, the estimated depth of the cloud is based on the 3D geometrical shape of the cloud and also is uncertain.
-Thus, we do not take into account the uncertainty of the gas volume density in the total uncertainty of the B-field strengths.
+Thus, we did not take into account the uncertainty of the gas volume density in the total uncertainty of the B-field strengths.
 For this reason, the uncertainty of might be underestimated.
-Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : .
-For the following analysis, we will employ only the map inferred from the polarimetric data at 154 m since it covers the larger area of the higher-density region in Oph-A than that from the data at 89 m.
 3.3 Measurement of magnetic field strength using other methods
@@
 This provides an improved measurement of B-field strength but also gives us the ratio of the turbulent to large-scale B-field strengths.
-Alternatively, Skalidis & Tassis (2021) introduced a modification of the DCF method, hereafter ST, assuming that the mean (ordered) and fluctuation components of the total B-field energy, and , respectively, are not independent.
-More recently, the differential measure approach (DMA; Lazarian et al., 2022) is considered promising to enhance the accuracy of B-field strength measurements at scales smaller than the turbulence injection scale and takes into account the anisotropic properties of MHD turbulence (given by factor ).
-For the sake of cross-checking the B-field strength measurement in Oph-A using the DCF method, we re-calculate the mean B-field strengths using ADF, ST, and DMA methods at the two subregions (a) and (e) associated with the densest parts of the cloud (these subregions are named as in Kwon et al.
-(2018), see also in Figure 15).
+Alternatively, Skalidis & Tassis (2021) introduced a modification of the DCF method (hereafter ST), assuming that the mean (ordered) and fluctuation components of the total B-field energy, and , respectively, are not independent.
+More recently, the differential measure approach (DMA; Lazarian et al., 2022) is considered promising to enhance the accuracy of B-field strength measurements at scales of smaller than the turbulence injection scale and takes into account the anisotropic properties of MHD turbulence (given by factor ).
+For the sake of cross-checking, the B-field strength measurement in Oph-A using the DCF method, we re-calculated the mean B-field strengths using the ADF, ST, and DMA methods at the two subregions (a) and (e) associated with the densest parts of the cloud (these subregions are named as in Kwon et al.
+(2018), see also in Fig.
+15).
 3.3.1 ADF method
@@
 where denotes an average, is the polarization angle difference between individual pairs of polarization vectors at position and , where and are the corresponding polarization angles at position and , respectively, and is number of pairs of vectors with a displacement of .
-At scales much smaller than the scale for a variation of the large-scale B-fields structure (typically of 5 seen by Planck observations), the total angular dispersion function can be expressed as:
+At scales much smaller than the scale for a variation of the large-scale B-field structure (typically of 5 seen by Planck observations), the total angular dispersion function can be expressed as:
 where represents the turbulent component of B-fields (), describes the large-scale structure of B-fields (), and is subject to measurement uncertainties.
 Also, the ratio between the turbulent and large-scale B-field strength indicates the angular dispersion, given as (Hildebrand et al., 2009):
-When , i.e., rad, one can obtain .
+When , i.e., rad, we can obtain .
 For each subregion (a) and (e) in Oph-A, we calculate the angular structure function using HAWC data at 154 m via Eq.
@@
 We select data points that 13.6″50″ to avoid fitting to structures higher than the beam size and to better fit to smaller-scale structures.
-Top panels in Figure 7 show structure functions of polarization angles as a function of the distance of displacement toward subregions (a) and (e).
+Top panels in Fig.
+7 show structure functions of polarization angles as a function of the distance of displacement toward subregions (a) and (e).
 From the fitted value of the fits, we derive values of and toward subregions (a) and (e), respectively.
-We find these values in good agreement with mean polarization angle dispersion values estimated using our approach described in Section 3.2.1, which are 9.2° and 14.3° in subregions (a) and (e), respectively.
+We find these values in good agreement with mean polarization angle dispersion values estimated using our approach described in Sect.
+3.2.1, which are 9.2° and 14.3° in subregions (a) and (e), respectively.
 Toward subregions (a) and (e), mean gas volume densities are and cm-3 respectively, and their mean non-thermal FWHMs are 1.74 and 1.45 km s-1, respectively.
 We substitute new values together with the above mean values of and into Eq.
-2 to re-calculate strengths in subregions (a) and (e), which are mG and mG, respectively.
+2 to calculate strengths in subregions (a) and (e), which are mG and mG, respectively.
 These values are in good agreement with strengths obtained using the DCF method, where values are in ranges of 0.32–1.55 mG and 0.15–0.97 mG toward subregions (a) and (e), respectively, with corresponding mean values of 0.84 and 0.37 mG.
@@
 Another advantage of the ST method is that this method accounts for the correlation between the fluctuation and mean components of the total B-field energy not to be zeros.
-Here, we measure the maps of strength () in band C and D using the formula as appeared in Skalidis & Tassis (2021):
-We show the maps of across Oph-A in band C and D in Figure 16.
+Here, we measure maps of strength () in band C and band D using the formula as appeared in Skalidis & Tassis (2021):
+We show the maps of across Oph-A in band C and band D in Fig.
+17.
 We find that value ranges from mG at 89 m and mG at 154 m, appeared to be lower than .
@@
 3.3.3 DMA method
-We also re-calculate trengths in two subregions (a) and (e) using the DMA method, , using Eq. 10 following Lazarian et al. (2022):
+We also calculate trengths in two subregions (a) and (e) using the DMA method, , using Eq. 10 following Lazarian et al. (2022):
 where and are the second order structure function of polarization angles and central velocity at an angular scale , is the correction factor depending on the anisotropic properties of MHD turbulence.
@@
 We calculate the structure function of the central velocity of HCO+(43) line toward subregions (a) and (e), using the similar form in Eq.
-6 but for the central velocity , shown in middle panels in Figure 7.
+6 but for the central velocity , shown in middle panels in Fig.
+7.
 We then measure strengths in two subregions (a) and (e) via Eq.
 10, using the information from the together with and the mean values of in the subregions.
-The as a function of displacement is shown in bottom panels of Figure 7 for subregions (a) and (e).
+The as a function of displacement is shown in bottom panels of Fig.
+7 for subregions (a) and (e).
 It can be seen that the resulting in the range is likely flatten within their uncertainty.
 Thus, in each subregion we take the mean of in this range to be the average value, which are 0.34 mG and 0.12 mG in subregions (a) and (e), respectively.
-These values are slightly higher than the resulting estimated using the ST method, but at the low-end of that measured using the DCF method, i.e., 2–3 times less than the mean values estimated using the DCF method presented in Section 3.2.
+These values are slightly higher than the resulting estimated using the ST method, but they are at the low-end of that measured using the DCF method (i.e., 2–3 times less than the mean values estimated using the DCF method presented in Sect.
+3.2).
 These differences are indeed expected since the DCF method somewhat overestimates the strength in the cloud.
-We will discuss this issue in Section 4.1.
+We discuss this issue further in Sect.
+4.1.
 3.4 Mass-to-flux ratio
-The relative importance of B-fields compared to gravity in the cloud can be quantified by the mass-to-magnetic-flux ratio (or so-called mass-to-flux ratio), , which can be estimated in units of the critical ratio, (Nakano & Nakamura, 1978): .
+The relative importance of B-fields compared to gravity in the cloud can be quantified by the mass-to-flux ratio, , which can be estimated in units of the critical ratio, (Nakano & Nakamura, 1978): .
 We measure in the entire region of Oph-A, using the formula introduced by Crutcher (2004):
 where and are in units of cm-2 and G, respectively.
-Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : .
-For the following analysis, we will employ only the map estimated using the DCF method and inferred from the polarimetric data in band D since it covers the larger area of the higher-density region in Oph-A than that from the data in band C.
-The top panel in Figure 8 shows the spatial distribution of the mass-to-flux ratio in Oph-A.
+Due to the lack of component measurements in the Oph-A cloud, we adopted the component as the total B-field strengths : .
+For the analysis, we employed only the map estimated using the DCF method and inferred from the polarimetric data in band D; this is because it covers a larger area of the higher density region in Oph-A than that coming from the data in band C.
+The top panel in Fig.
+8 shows the spatial distribution of the mass-to-flux ratio in Oph-A.
 It reveals an increasing gradient of from the central and eastern parts of the clump to the outer regions, spanning from 0.03–2.56, with a median value of 0.26.
-The region exhibiting the lowest values of 1 is close to the OphS1 star, suggesting that the region is firmly magnetically sub-critical, i.e., B-fields are strong enough to prevent the gravitational collapse.
+The region exhibiting the lowest values of 1 is close to the OphS1 star, suggesting that the region is firmly magnetically sub-critical, namely, B-fields are strong enough to prevent the gravitational collapse.
 This can also be explained by the influence of the strong winds and radiation fields from the OphS1 star which tend to push the material away to to its vicinity at the west side.
 The mass-to-flux ratio increases toward the outskirt regions of the Oph-S1 star, especially toward regions along the ridge of the dense cloud (i.e.
-center, north, and south-east parts of the map) where the gas column density is cm-2.
+center, north, and southeast parts of the map) where the gas column density is cm-2.
 These regions are magnetically super-critical and so, the self-gravitational collapse might have occurred there, resulting in several protostars and a single protostellar core identified in the region (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
@@
 The factor in Eq.
-12 assumes that turbulence is isotropic, and allows the calculation of the three-dimensional (3D) Alfvénic Mach numbers.
+12 assumes that turbulence is isotropic, and allows the calculation of the 3D Alfvénic Mach numbers.
 We note that adding this factor will provide upper limits to the estimated .
-The middle panel in Figure 8 shows the map of toward Oph-A using results of polarimetric data at 154 m and HCO+(4–3) as the gas tracer.
+The middle panel in Fig.
+8 shows the map of toward Oph-A using results of polarimetric data at 154 m and HCO+(4–3) as the gas tracer.
 The values are in the range from 0.1–1.5 across the cloud, with a median value of 0.38.
 Most of Oph-A shows ¡1 (magnetically sub-Alfvénic), suggesting the B-field dominates turbulence in the cloud.
-However, two small regions at the north and southwest of the map exhibit 1, indicating strong Alfvénic motions of the gas (magnetically supper-Alfvénic).
+However, two small regions at the north and southwest of the map exhibit 1, indicating strong Alfvénic motions of the gas (magnetically super-Alfvénic).
 These regions are also associated with the positions of a few prestellar and protostellar cores, and YSO candidates (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
 3.6 Plasma parameter
-To compare the magnetic presure () and thermal pressure () in Oph-A, we construct the map of the plasma beta parameter (), given as:
-where is the thermal sound speed measured in units of km s-1, assuming = (see Section 3.2.2).
-The bottom panel in Figure 8 shows the map of the parameter toward Oph-A in band D.
-We obtain ¡1 across the map with a median value of 0.02, implying that the cloud is supported by strong magnetic pressure which is much more prominent than the one induced by the thermal motions of the gas.
+To compare the magnetic presure () and thermal pressure () in Oph-A, we constructed the map of the plasma beta parameter (), given as:
+where is the thermal sound speed measured in units of km s-1, assuming = (see Sect. 3.2.2).
+The bottom panel in Fig.
+8 shows the map of the parameter toward Oph-A in band D.
+We obtained ¡1 across the map with a median value of 0.02.
+This result implies that the cloud is supported by strong magnetic pressure, which is much more prominent than that induced by the thermal motions of the gas.
 3.7 Virial mass analysis
-To probe the dynamical state of the Oph-A cloud, we calculate the virial parameter (), which is the ratio of the virial mass () and the isothermal mass () of the cloud.
+To probe the dynamical state of the Oph-A cloud, we calculated the virial parameter (), which is the ratio of the virial mass () and the isothermal mass () of the cloud.
 Here, we assume the cloud as a uniform density sphere with an effective radius of as a whole, having a mean mass density of .
@@
 The virial mass, , determined from magnetic, thermal, and non-thermal kinetic energies can be estimated as (see e.g., Bertoldi & McKee, 1992; Pillai et al., 2011):
-where , , and are respectively the mean values of the Alfvénic velocity, thermal sound speed, and of non-thermal velocity dispersion toward the cloud, is the gravitational constant.
-We calculate by selecting an ellipse covering the area where is measured.
+where , , and are the mean values of the Alfvénic velocity, thermal sound speed, and of non-thermal velocity dispersion toward the cloud, respectively, while is the gravitational constant.
+We calculate by selecting an ellipse covering the area where was measured.
 This ellipse is centered at (RA, Dec) = () with the semi-major and semi-minor axes are =120 and =75, respectively.
@@
 The corresponding mean values of velocities are: , , and .
-We use these mean values of the velocities, non-thermal velocity dispersion, and the gas volume density to calculate and in Eqs.
+We used these mean values of the velocities, non-thermal velocity dispersion, and the gas volume density to calculate and in Eqs.
 14 and 15, yielding 15.
 This shows that the cloud is gravitationally unbound and that B-fields and turbulence can support the cloud against self-gravity.
-This high value of obtained in the cloud overalls could explain the low efficiency of star formation in rho Oph-A though in a few regions the density is very high up to 105 or 106 cm-3.
-This is consistent with the lack of prestellar cores and protostellar sources in the central region, which is seen in the lower-density parts of the cloud only (see e.g., Figure 8).
-In order to study smaller scales across the clump, polarimetric observations with a higher spatial resolution would be required.
+This high value of obtained in the cloud overalls could explain the low efficiency of star formation in Oph-A; however, in a few regions, the density is very high (up to 105 or 106 cm-3).
+This is consistent with the lack of prestellar cores and protostellar sources in the central region, which can be seen in the lower density parts of the cloud only (see e.g., Fig.
+8).
+To study smaller scales across the clump, polarimetric observations with a higher spatial resolution would be required.
 4 Discussion
 4.1 Morphology and strengths of magnetic field in Oph-A
-The B-field orientation in Oph-A is well-ordered and predominantly perpendicular to the ridge of the dense cloud (Section 3.1).
+The B-field orientation in Oph-A is well-ordered and predominantly perpendicular to the ridge of the dense cloud (Sect.
+3.1).
 The feature is illustrated most clearly by the data at the longer wavelength (154 m).
-The data at 89 m better traces warmer dust in the lower-density region of Oph-A, which surrounds the Oph-S1 star.
-The B-field properties measured using SOFIA/HAWC polarization data are qualitatively consistent with previous measurements of B-field orientations in Oph-A obtained at 850 m using SCUBA-2/POL–2 (Kwon et al., 2018) and SCUBA/SCUPOL (Matthews et al., 2009) on JCMT and at NIR wavelengths (JHK bands) using IRSF/SIRPOL (Kwon et al., 2015).
-In the POL–2 dataset, Oph-A was divided into 10 subregions, identified based on their B-field position angle and the polarization fraction, covering the entire extent of HAWC maps (see Figure 15).
+The data at 89 m better traces warmer dust in the lower density region of Oph-A, which surrounds the Oph-S1 star.
+The B-field properties measured using SOFIA/HAWC polarization data are qualitatively consistent with previous measurements of B-field orientations in Oph-A obtained at 850 m using SCUBA-2/POL-2 (Kwon et al., 2018) and SCUBA/SCUPOL (Matthews et al., 2009) on JCMT and at NIR wavelengths (JHK bands) using IRSF/SIRPOL (Kwon et al., 2015).
+In the POL-2 dataset, Oph-A was divided into ten subregions, identified based on their B-field position angle and the polarization fraction, covering the entire extent of HAWC maps (see Fig.
+15).
 The highest density region is characterized by an average B-field position angle of 50°, which is also consistent with the NIR data toward the L1688 main cloud (Kwon et al., 2015).
-Table 1 shows the median B-field position angle () in each subregion which has been covered also by the FIR and sub-mm observations.
+Table 1 shows the median B-field position angle () in each subregion, which has been covered by the FIR and sub-mm observations a well.
 The median B-field position angles from the 89, 154, and 850 m datasets are in rather good agreement, with the difference in values ranging from 3 to 19 (30%).
@@
 | (g) | 25 | 14 | 0.5 | 0.7 | – | |||
-(i) | 23 | 13 | 0.4 | 0.5 | – |
+| (i) | 23 | 13 | 0.4 | 0.5 | – | |||
 Notes: (a) The subregions are named as in Kwon et al. (2018). (b) The relative difference in is taken as , where is the B-field position angle at wavelength =[89, 154] (m).
@@
 (2018) calculated in three subregions associated with the densest parts of Oph-A using non-thermal velocity dispersion from the N2H+(1–0) line (André et al., 2007) and the sub-mm data at 850 m.
-In our study, we calculate pixel-by-pixel toward the entire Oph-A region.
-To facilitate comparisons with results obtained from the sub-mm data, we calculate the mean value of in two out of these three subregions, covered by the HAWC data (see Table 1).
+In our study, we calculated the pixel-by-pixel toward the entire Oph-A region.
+To facilitate comparisons with results obtained from the sub-mm data, we calculated the mean value of in two out of these three subregions, covered by the HAWC data (see Table 1).
 Mean values of using FIR data are lower than that obtained with the sub-mm data, which traces colder dust layers in the cloud characterized also by the higher densities.
-The HAWC data, on the other hand, essentially traces the warmer layers in the lower-density region.
-Also, Kwon et al.
-(2018) adopted values of the polarization angle dispersion as the standard deviation of the polarization angles, which are much smaller than those using the spread of the dust polarization angles (see Section 3.2.1).
+The HAWC data, on the other hand, essentially traces the warmer layers in the lower density region.
+In addition, Kwon et al.
+(2018) adopted values of the polarization angle dispersion as the standard deviation of the polarization angles, which are much smaller than those using the spread of the dust polarization angles (see Sect.
+3.2.1).
 As a consequence, the obtained with sub-mm data is expected to be stronger than the one obtained with FIR data.
@@
 The mass-to-flux ratio, Mach Alfvénic number, and plasma parameter will decrease, but this will not significantly change our conclusion that B-fields are generally dominant over gravity, turbulent, and gas kinetic energies, especially toward the center of the cloud.
-Furthermore, as presented in Section 3.3, we test that the strengths obtained with the DCF method in subregions (a) and (e) are rather in good agreement with values obtained with the ADF method but higher than those using the ST and DMA methods (see Table 2).
+Furthermore, as presented in Sect.
+3.3, we test that the strengths obtained with the DCF method in subregions (a) and (e) are rather in good agreement with values obtained with the ADF method but higher than those using the ST and DMA methods (see Table 2).
 We use new values estimated with ADF, ST, and DMA methods toward two subregions (a) and (e) to estimate the mass-to-flux ratio, Mach Alfvénic number, and plasma, which are all higher than that measured with the DCF method (see Table 2).
@@
 A similar behavior is seen in subregion (e) except that this subregion changes from trans- to super-Alfvénic (see Table 2).
-We now assume that the B-field strength in the entire cloud decreases by a factor of 4.5 as measured with the ST method.
-In this case, the majority of the cloud will become magnetically trans/super-critical (1), except for the regions in the vicinity of the Oph–S1 star (see the top panel of Figure 20).
+We go on to assume that the B-field strength in the entire cloud decreases by a factor of 4.5, as measured with the ST method.
+In this case, the majority of the cloud will become magnetically trans- or super-critical (1), except for the regions in the vicinity of the Oph–S1 star (see the top panel of Fig.
+20).
 However, this trend of remains similar as obtained using .
-The cloud will become mostly super-Afvénic with ¿1, except for the southeast region of the cloud (see the middle panel of Figure 20).
+The cloud will become mostly super-Afvénic with ¿1, except for the southeast region of the cloud (see the middle panel of Fig.
+20).
 This assumption will also result in the Alfvénic velocity decreasing by a factor of 4.5 and thus, .
@@
 To investigate what is the relative orientation of the B-fields with respect to the gas structure in Oph-A, we quantify the projected offset angle between the orientations of B-field vectors projected in the plane-of-the-sky and the main structure associated with the ridge of the cloud, given as , where is the position angle of the ridge.
-We identify the shape of the ridge using the Python package RadFil444https://github.com/catherinezucker/radfil (Zucker & Chen, 2018, see appendix F).
+We identified the shape of the ridge using the Python package RadFil444https://github.com/catherinezucker/radfil (Zucker & Chen, 2018, see Appendix E).
 Figure 9 shows the map of offset angles () in Oph-A, overlaid by contours of continuum emission at 154 m.
-Clearly, the offset angle appears to increase when moving from the lower-density regions to the denser regions, indicating that the relative orientation of B-fields with respect to the main structure of the cloud turns from parallel to perpendicular toward the high-density region.
-This is in agreement with Planck polarization observations at 353 GHz (850 m) toward the entire L1688 region where the relative orientation of B-fields with respect to the gas structure changes gradually from parallel in the low-density region to perpendicular toward the higher-density region, with the transition occurring at cm-2 corresponding to the visual extinction of 3.3 mag555We convert from using the relation: molecules cm-2 mag-1 with a “standard” assumption of the total-to-selective extinction ratio, , of 3.1 (Frerking et al., 1982).
+Clearly, the offset angle appears to increase when moving from the lower density regions to the denser regions, indicating that the relative orientation of B-fields with respect to the main structure of the cloud turns from parallel to perpendicular toward the high-density region.
+This is in agreement with Planck polarization observations at 353 GHz (850 m) toward the entire L1688 region where the relative orientation of B-fields, with respect to the gas structure changes gradually.
+It runs parallel in the low-density region to perpendicular toward the higher density region, with the transition occurring at cm-2 and corresponding to the visual extinction of 3.3 mag555We convert from using the relation: molecules cm-2 mag-1 with a “standard” assumption of the total-to-selective extinction ratio, , of 3.1 (Frerking et al., 1982).
 (Soler, 2019).
-This transition is also observed in other nearby molecular clouds in the Gould Belt using Planck 353 GHz polarization observations, typically occurring at cm-2 or 2.7 mag (see e.g., Planck Collaboration et al., 2016; Soler et al., 2017).
+This transition has also been observed in other nearby molecular clouds in the Gould Belt using Planck 353 GHz polarization observations, typically occurring at cm-2 or 2.7 mag (see e.g., Planck Collaboration et al., 2016; Soler et al., 2017).
 This phenomenon can be explained by the motion of the high-density material during the self-gravitational collapse.
@@
 Eventually, the B-field lines are dragged and reshaped to become more uniform and perpendicular to the high-density structure of the field.
-The HAWC data traces B-fields in the dense region of Oph-A with cm-2 (see Figure 1), which corresponds to in range of 3.2–256 mag.
-Thus, we further investigate in detail the relation between the relative orientation of B-fields with respect to the cloud structure and the within Oph-A, as shown in Figure 10.
+The HAWC data traces B-fields in the dense region of Oph-A with cm-2 (see Fig.
+1), which corresponds to in range of 3.2–256 mag.
+Thus, we further investigate in detail the relation between the relative orientation of B-fields with respect to the cloud structure and the within Oph-A, as shown in Fig.
+10.
 Clearly, in the regime where 30 mag the offset angle () decreases with and drops from 60° at =4 mag down to 0° at 20 mag.
 This indicates that the relative orientation of B-fields changes from perpendicular back to parallel at 18–33 mag.
-Similar trends of this transition were found in the dense region of the Serpens South filament (Pillai et al., 2020) and in Serpens Main dense cloud (Kwon et al., 2022), occurring at 21 mag and 50 mag, respectively.
+Similar trends of this transition were found in the dense region of the Serpens South filament (Pillai et al., 2020) and in the Serpens Main dense cloud (Kwon et al., 2022), occurring at 21 mag and 50 mag, respectively.
 This is likely because the B-fields are already dragged into the flow of very dense material during the gravitational contraction.
 Interestingly, we observe that the offset angle then progressively increases with 40 mag in our region, indicating the relative orientation of B-field with respect to the ridge changes back from parallel to perpendicular (with the transition at 42–60 mag).
-We note, however, that this feature was not seen in the Serpens South filamentary cloud whose maximum 60 mag (see Figure 3 in Pillai et al.
-(2020)) but in the Serpens Main cloud region where 170 mag (see Figure 6 in Kwon et al.
+We note, however, that this feature was not seen in the Serpens South filamentary cloud, whose maximum is 60 mag (see Fig.
+3 in Pillai et al.
+(2020)); however, it was seen in the Serpens Main cloud region, where 170 mag (see Fig.
+6 in Kwon et al.
 (2022)).
@@
 The B-field strength ( or ) can be expressed as a function of (): .
-Theoretical models predict 2/3 in the weak B-field regions (Mestel, 1966) and 0.5 in the strong B-field regions where magnetic energy dominates over the gravity (Mouschovias & Ciolek, 1999).
+Theoretical models predict 2/3 in the weak B-field regions (Mestel, 1966) and 0.5 in the strong B-field regions, where magnetic energy dominates over the gravity (Mouschovias & Ciolek, 1999).
 Consequently, in the latter case, it suggests that would increase faster than if there is any collapse in the dense cloud.
-Here, we show in Figure 11 the relation toward the Oph-A cloud, using B-fields measured with the DCF method in band D.
+Here, we show in Fig.
+11 the relation toward the Oph-A cloud, using B-fields measured with the DCF method in band D.
 We find that for densities , the B-fields distribution tends to be flattened.
 In a regime with densities higher than we start seeing the increasing trend of B-fields with respect to gas densities.
-Thus, we decide to test the relation at the densest part of the cloud only.
-We choose this region as the area within the third contour of Stokes- at band D (see in the bottom right panel of Figure 5), where Stokes.
+Thus, we decided to test the relation at the densest part of the cloud only.
+We chose this region as the area within the third contour of Stokes- at band D (see in the bottom right panel of Fig.
+5), where Stokes.
 Within this region, it reveals a relatively weak correlation between the B-fields and densities (with the Pearson coefficient of the correlation is found , corresponding to a significance of ).
-We perform the power-law fit using the Python curve-fit routine (), using mean values of estimated within every interval of =0.08666The bin width is chosen by using the Freedman-Diaconis rule.
-, which results in a power-law index of .
-In Figure 11, we also overplot the results of using other methods in subregions (a) and (e).
-We find that the fitted line well covers the data point representing the resulting value using ADF method and shows good agreement with those estimated with ST and DMA in subregion (a) within of uncertainty.
+We performed the power-law fit using the Python curve-fit routine (), using mean values of estimated within every interval of =0.08666The bin width is chosen by using the Freedman-Diaconis rule.
+.
+This gives a power-law index of .
+In Fig.
+11, we also overplot the results of using other methods in subregions (a) and (e).
+We find that the fitted line well covers the data point representing the resulting value using ADF method and shows good agreement with those estimated with ST and DMA in subregion (a) within and uncertainty of .
 The resulting is found consistent with theoretical predictions for strong B-fields (Mouschovias & Ciolek, 1999).
-This is also consistent with , representing strong B-fields in Oph-A (see bottom panel in Figure 8 and Section 3.6).
-We further perform the fit to values measured with the ST method from Sect.
+This is also consistent with , representing strong B-fields in Oph-A (see bottom panel in Fig.
+8 and Sect.
+3.6).
+We then performed the fit to values measured with the ST method from Sect.
 3.3.2 and find .
-We find a similar power-law index for the dataset at band C, which are and using result of and , respectively (see Appendix E).
+We found a similar power-law index for the dataset at band C, namely, and using result of and , respectively (see Appendix C).
 4.4 Magnetic fields versus stellar feedback
@@
 The radiation pressure acting on both dust and gas (assuming dust is well-coupled to the gas in the cloud) can drive the material away into the denser region.
-The lower-density matter in the outer regions of Oph-A is more exposed to the ionization front and pushed away faster than the higher-density matter in the inner regions.
+The lower density matter in the outer regions of Oph-A is more exposed to the ionization front and pushed away faster than the higher density matter in the inner regions.
 This results in dust and gas being compressed and forming the dense shell structure at the edge of the cloud.
@@
 Figure 12 shows the map of the continuum emission at 154 m from SOFIA/HAWC+, superimposed by the white contours showing the H2 column density.
-We calculate the external radiation pressure of the Oph-S1 star acting on dust and gas in the densest part of the cloud, given by:
-where =1100 is the luminosity of the Oph-S1 star adopted from Lada & Wilking (1984) and 0.06 pc is the projected distance to the Oph-S1 star.
-We obtain K cm-3, which is in agreement with the result calculated from the ionizing photons emitted by the Oph-S1 star (2 K cm-3, Pattle et al., 2015).
+We calculated the external radiation pressure of the Oph-S1 star acting on dust and gas in the densest part of the cloud, given by:
+where =1100 is the bolometric luminosity of the Oph-S1 star adopted from Lada & Wilking (1984) and 0.06 pc is the projected distance from the cloud to the Oph-S1 star.
+We obtained K cm-3, which is in agreement with the result calculated from the ionizing photons emitted by the Oph-S1 star (2 K cm-3, Pattle et al., 2015).
 Here, we omit the fact that the UV radiation from the Oph-S1 star can be absorbed by intervening dust in between the star and the Oph-A region.
@@
 16.
-To compare with the external radiation pressure, we calculate the internal thermal pressure due to the hot gas in Oph-A, .
-We consider the hot gas within the densest part of Oph-A where ¿0.4 Jy/arcsec2 at 154 m only, which corresponds to the area inside the third contour shown in bottom right panel of Figure 5.
-Within this region, we adopt in the range of 104–106 cm-3 and the average temperature of the hot gas 20 K, which yields in the range of (0.2–20) K cm-3.
-With this assumption, we note that could be underestimated since there is a gas heating process in the cloud dominated by the UV radiation from the Oph-S1 star and by the photoelectric emission from dust exposed to the UV radiation (i.e., ).
+To make a comparison with the external radiation pressure, we calculated the internal thermal pressure due to the hot gas in Oph-A, .
+We considered the hot gas within the densest part of Oph-A, where Stokes-0.4 Jy/arcsec2 at 154 m only, which corresponds to the area inside the third contour shown in bottom right panel of Fig.
+5.
+Within this region, we adopted in the range of 104–106 cm-3 and the average temperature of the hot gas 20 K, which yields in the range of (0.2–20) K cm-3.
+With this assumption, we note that could be underestimated, since there is a gas heating process in the cloud dominated by the UV radiation from the Oph-S1 star and by the photoelectric emission from dust exposed to the UV radiation (i.e., ).
 The highest internal thermal pressure of 20106 K cm-3 is seen toward the densest part of the cloud, associated with the position of the SM1 core.
 This suggests that the internal thermal pressure is strong enough to help the material refrain from external pressure.
-We note that, however, in the vicinity region away from the center (for e.g., the tails of the ridge shown in Figure 12), the internal thermal pressure is weaker, at least one order of magnitude less than the external one.
+We note that, however, in the vicinity region away from the center (e.g., the tails of the ridge shown in Fig.
+12), the internal thermal pressure is weaker, at least one order of magnitude less than the external one.
 In this case, the radiation pressure from the Oph-S1 star on the east side is dominant and able to push away all the matter into the surrounding environment toward the west side.
-However, it is clear that the material in the low-density region still survives, despite owning a low internal thermal pressure (see Figure 12), indicating that there must be another mechanism at play, for example, the presence of strong B-fields of a few mG in the cloud.
-We estimate roughly the magnetic pressure () to be 107–108 K cm-3 for B-field strengths between 1–2 mG.
+However, it is clear that the material in the low-density region still survives, despite owning a low internal thermal pressure (see Fig.
+12).
+This indicates that there must be another mechanism at play, for example, the presence of strong B-fields of a few mG in the cloud.
+We roughly estimated the magnetic pressure () to be 107–108 K cm-3 for B-field strengths between 1–2 mG.
 This indicates that the magnetic pressure is at least two orders of magnitude higher than the external pressure of 106 K cm-3, sufficient to support the matter against the external pressure along with the gas thermal pressure.
-We further investigate the interplay between B-fields and radiation directions on the evolution of the cloud.
-We determine the main radiation direction from the Oph-S1 star toward the densest region associated with the position of the starless core SM1 as the sky-projected direction from the Oph-S1 star and SM1 core.
-We find that the main radiation direction is 70°, measured east of north.
-The main B-field orientation toward the densest region of Oph-A of 74°, almost parallel to the radiation field.
+We further investigated the interplay between B-fields and radiation directions on the evolution of the cloud.
+We determined the main radiation direction from the Oph-S1 star toward the densest region associated with the position of the starless core SM1 as the sky-projected direction from the Oph-S1 star and SM1 core.
+We found that the main radiation direction is 70°, measured east of north.
+The main B-field orientation toward the densest region of Oph-A of 74° is almost parallel to the radiation field.
 The strong radiation pressure from the Oph-S1 star pushes all the material away and creates a dense shell toward the west side.
@@
 (2009).
-However, in the northeast and southeast regions the offset angles between the B-field vectors and radiation field are larger, i.e., the B-field orientation becomes perpendicular to the radiation direction.
-In this case, the magnetic pressure and tension induced by B-fields run parallel with the radiation pressure, suggesting magnetic pressure can support the matter to confront the radiative compression and prevent the sweeping dust and gas away.
+However, in the northeast and southeast regions, the offset angles between the B-field vectors and radiation field are larger; namely, the B-field orientation becomes perpendicular to the radiation direction.
+In this case, the magnetic pressure and tension induced by B-fields run parallel with the radiation pressure, suggesting magnetic pressure can support the matter to confront the radiative compression and prevent the dust and gas from being swept away.
 This results in bending the material structure and forming a dense ridge with an umbrella-like shape.
-Interestingly, in the case study of Oph-A the ionization source is very close to the cloud (¡0.1 pc) and the offset position angle between the radiation and B-fields is not homogeneous but varies from parallel to orthogonal.
-This allows us to examine widely how the interplay between the B-fields and radiation fields could influence the material structure of the cloud.
+Interestingly, in the case study of Oph-A the ionization source is very close to the cloud (¡0.1 pc) and the offset position angle between the radiation and B-fields is not homogeneous; instead, it varies from parallel to orthogonal.
+This allows us to closely examine how the interplay between the B-fields and radiation fields could influence the material structure of the cloud.
 5 Summary
-In this paper, we presented the analysis using the SOFIA/HAWC polarimetric observations toward the Oph-A molecular cloud in the L1688 active star-forming region of the Ophiuchus dark cloud complex.
+In this paper, we present our analysis of the Ophiuchus dark cloud complex, using the SOFIA/HAWC polarimetric observations toward the Oph-A molecular cloud in the L1688 active star-forming region.
 Our main results are as follows:
 Using the dust polarization maps at 89 and 154 m, we derived the map of B-field morphology in Oph-A.
-We found that B-fields are generally well-ordered and primarily perpendicular to the ridge of the cloud toward the higher-density regime.
+We found that B-fields are generally well-ordered and primarily perpendicular to the ridge of the cloud toward the higher density regime.
 We estimated in the entire region of Oph-A using the DCF method, ranging from 0.2 to 2.3 mG at 89 and to 2.5 mG at 154 m.
-We found is prominent toward the densest region of the cloud associated with the starless core SM1 and decreases toward the outskirts.
-This is the first time we estimate the distribution of the B-field strengths in Oph-A, which is essential to understanding how B-fields vary in the molecular cloud globally.
-Our derived B-fields morphology and strengths well agree with previous measurements using the sub-mm dust polarization (Kwon et al., 2018).
-We also estimated strength using more advanced techniques, including the ADF, ST, and DMA, toward two subregions associated with the densest part of the cloud.
-While strengths estimated with the ADF method were in agreement with those from DCF, the ST and DMA methods provided lower values of 4.5 and 2-3 times, respectively.
-We studied the role of B-fields relative to gravity and turbulence by calculating the maps of the mass-to-flux ratio, Alfvénic Mach number, and plasma parameter toward Oph-A using the inferred from SOFIA/HAWC data at 154 m.
+We found that is prominent toward the densest region of the cloud associated with the starless core SM1 and decreases toward the outskirts.
+This is the first time we have successfully estimated the distribution of the B-field strengths in Oph-A, which is essential to understanding how B-fields vary in the molecular cloud globally.
+Our derived B-field morphology and strengths well agree with previous measurements using the sub-mm dust polarization (Kwon et al., 2018).
+We also estimated the strength using more advanced techniques, including the ADF, ST, and DMA, toward two subregions associated with the densest part of the cloud.
+While the strengths estimated with the ADF method were in agreement with those from DCF, the ST and DMA methods provided lower values of 4.5 and 2-3 times, respectively.
+We studied the role of B-fields relative to gravity and turbulence by calculating the maps of the mass-to-flux ratio, Alfvénic Mach number, and plasma parameter toward Oph-A, using the inferred from SOFIA/HAWC data at 154 m.
 We found that toward the central dense parts the cloud is magnetically sub-critical and sub-Alfvénic where B-fields dominate over the gravity and turbulence.
-We obtained in Oph-A, suggesting the cloud is supported by the strong magnetic pressure and much greater than the thermal gas pressure.
+We obtained a result of in Oph-A, suggesting the cloud is supported by the strong magnetic pressure and much greater than the thermal gas pressure.
 The lower strengths obtained with the ST and DMA methods do not significantly change our conclusions on the roles of B-fields relative to gravity and turbulence on star formation.
@@
 In contrast, Oph-A is known as one of the densest parts of the active star-forming region L1688 hosting a few pre- and proto-stellar cores and embedded protostars.
-Higher spatial resolution observations toward the smaller scale region in the cloud will be necessary to investigate in detail the dynamical state of the cloud.
+Higher spatial resolution observations toward the smaller scale region in the cloud will be necessary to investigate the dynamical state of the cloud in greater detail.
 We examined the role of B-fields in supporting the cloud against the radiation pressure from the nearby high-mass star Oph-S1.
-We found that at the center part of the cloud, the magnetic pressure indeed does not play a role in helping the material prevent radiation pressure.
-However, toward the northeast and southeast regions of the cloud, the magnetic pressure positively supports the cloud to prevent the radiative pressure together with the gas thermal pressure.
+We found that at the center part of the cloud, the magnetic pressure indeed does not play a role in helping the material to avoid the onset of radiation pressure.
+However, toward the northeast and southeast regions of the cloud, the magnetic pressure positively supports the cloud to prevent both the effects of radiative pressure and gas thermal pressure.
 Acknowledgements.
@@
 N.L.
-thanks Thang Nguyen Tat for his valuable comments as part of the 3D structure of the Oph-A.
+thanks Nguyen Tat Thang for his valuable comments as part of the 3D structure of the Oph-A.
 N.L.
@@
 was funded by the Master, Ph.D.
-Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2022.TS083.
+Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2023.TS.077.
 M.H.
@@
 also acknowledges support by the grant from the Simons Foundation to IFIRSE, ICISE (916424, PI: N.H.).
-P.N.D., N.B.N., and N.T.P.
+P.N.D.
+and N.T.P.
 were funded by the Vietnam Academy of Science and Technology under project code THTETN.03/24-25.
@@
 - Hoang et al. (2022) Hoang, T., Tram, L. N., Minh Phan, V. H., et al. 2022, AJ, 164, 248
+- Holland et al. (2013) Holland, W. S., Bintley, D., Chapin, E. L., et al. 2013, MNRAS, 430, 2513
 - Hwang et al. (2021) Hwang, J., Kim, J., Pattle, K., et al. 2021, ApJ, 913, 85
 - Hwang et al. (2023) Hwang, J., Pattle, K., Parsons, H., Go, M., & Kim, J. 2023, AJ, 165, 198
+- Kandori et al.
+(2006) Kandori, R., Kusakabe, N., Tamura, M., et al.
+2006, in Ground-based and Airborne Instrumentation for Astronomy, ed.
+I.
+S.
+McLean & M.
+Iye, Vol.
+6269, International Society for Optics and Photonics (SPIE), 626951
 - Kauffmann et al. (2008) Kauffmann, J., Bertoldi, F., Bourke, T. L., Evans, N. J., I., & Lee, C. W. 2008, A&A, 487, 993
@@
 Appendix A Histogram of SOFIA/HAWC dataset
-Figures 13 and 14 show the histogram of Stokes-, , SNRI, , , and SNRp of the SOFIA/HAWC data in band C and band D.
+Figures 13 and 14 show histograms of Stokes-, , SNRI, , , and SNRp of the SOFIA/HAWC data in band C and band D.
 Table 3 shows the summary of the mean and RMS of the distribution of Stokes-, , their uncertainties and SNRs.
@@
 | Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | Mean | RMS | |
-89 m | 0.163 | 0.166 | 0.005 | 0.004 | 44.8 | 52.9 | 8.3 | 9.2 | 6.3 | 7.8 | 3.2 | 3.5 |
-154 m | 0.143 | 0.142 | 0.001 | 0.001 | 177.5 | 214.1 | 5.7 | 4.1 | 1.2 | 1.5 | 10.3 | 9.9 |
+| 89 m | 0.163 | 0.166 | 0.005 | 0.004 | 44.8 | 52.9 | 8.3 | 9.2 | 6.3 | 7.8 | 3.2 | 3.5 |
+| 154 m | 0.143 | 0.142 | 0.001 | 0.001 | 177.5 | 214.1 | 5.7 | 4.1 | 1.2 | 1.5 | 10.3 | 9.9 |
 Appendix B Magnetic field orientations and strengths at different wavelengths
-In Figure 15, we compare the B-fields vectors obtained from HAWC data at 89 and 154 m (this study), and from POL-2 data at 850 m (Kwon et al.
+In Fig.
+15, we compare the B-fields vectors obtained from HAWC data at 89 and 154 m (this study), and from POL-2 data at 850 m (Kwon et al.
 2018).
 Vectors show the mean B-field position angles at the subregions in Oph-A.
-Figure 16 shows the maps of B-field strength estimated using the ST method () at 89 and 154 m. Figure 17 shows the histogram distribution of the obtained with HAWC+ data at 89 and 154 m.
+Figure 16 shows the histogram distribution of the estimated obtained with HAWC data at 89 and 154 m. Figure 17 shows the maps of B-field strength estimated using the ST method () at 89 and 154 m.
 Appendix C Velocity dispersion traced by other molecular lines
@@
 To each pixel we fit the spectrum with the sum of 3 single Gaussian profiles and then use the line width of the primary peak as the toward the computing pixel.
-We construct the map of the FWHM velocity dispersion, following the procedure presented in Section 3.2.2 for the case of the HCO+ molecular line.
+We construct the map of the FWHM velocity dispersion, following the procedure presented in Sect.
+3.2.2 for the case of the HCO+(43) molecular line.
 Figure 18 shows the map of using the N2H+(32) molecular line toward Oph-A.
 The values ranges from 0.18 to 1.73 km s-1 with a median value of 0.95 km s-1, in good agreement with those obtained by HCO+(43) line.
-We note, however, that the spatial extent of the HCO+(43) emission line is large than that of the N2H+(32).
+We note, however, that the spatial extent of the HCO+(43) emission line is largeer than that of the N2H+(32).
 Thus, we use the map of FWHM of the HCO+(43) line to calculate the strengths in Oph-A.
@@
 (2007) estimated an average value of the 1D nonthermal velocity dispersion toward Oph-A, =0.2 km s-1, corresponding to 0.47 km s-1.
-Observations of the NH3 hyperfine emission from the Green Bank Telescope (GBT) also returned a similar 1D velocity dispersion with the average value of 0.3 km s-1, corresponding to an average of FWHM of 0.65 km s-1 (Chen et al.
+Observations of the NH3 hyperfine emission from the Green Bank Telescope also returned a similar 1D velocity dispersion with the average value of 0.3 km s-1, corresponding to an average of FWHM of 0.65 km s-1 (Chen et al.
 2019).
-Appendix D Mass-to-flux ratio, Mach Alfvén Mach number, and beta plasma
-Figure 19 show the maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C.
-Figure 20 show maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C (left panels) and D (right panels).
-Appendix E relation using estimated with the ST method
-Figure 21 shows the relation between and where B-field strength is estimated using the St method.
-Appendix F Main structure of the ridge identification
-We identify the elongated shape of the ridge in the densest region of Oph-A using the Python package RadFil.
-We first apply a threshold of 0.5 Jy/arcsec2 in the map of Stokes- at 154 m to build a mask for the main shape of the densest region (the ridge).
-The spine of the ridge is identified by using the Python package FilFinder777https://github.com/e-koch/FilFinder from Koch & Rosolowsky (2015).
-We then smooth the ridge spine and define a tangent line at every position on the ridge where the orientation of the ridge () at every position is identified.
+Appendix D Mass-to-flux ratio, Alfvénic Mach number, beta plasma, and relation results obtained by using
+Figure 19 shows the maps of mass-to-flux ratio (top panel), Alfvénic Mach number (middle panel), and beta plasma (bottom panel) toward Oph-A using in band C.
+Figure 20 show maps of mass-to-flux ratio (top), Alfvénic Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C (left panels) and D (right panels).
+Figure 21 shows the relation between and using in band D.
+Appendix E Main structure of the ridge identification
+We identified the elongated shape of the ridge in the densest region of Oph-A using the Python package RadFil.
+We first applied a threshold of Stokes-0.5 Jy/arcsec2 in the map of Stokes- at 154 m to build a mask for the main shape of the densest region (the ridge).
+The spine of the ridge was identified by using the Python package FilFinder777https://github.com/e-koch/FilFinder from Koch & Rosolowsky (2015).
+We then smoothed the ridge spine and define a tangent line at every position on the ridge, where the orientation of the ridge () at every position was identified.
 Figure 22 shows the schematic view of the ridge and the perpendicular cuts at each position on the ridge.